### v1

# NAMO-LLM: Efficient Navigation Among Movable Obstacles with Large Language Model Guidance

Abstract
Several planners have been proposed to compute robot paths that reach desired goal regions while avoiding obstacles. However, these methods fail when all pathways to the goal are blocked. In such cases, the robot must reason about how to reconfigure the environment to access task-relevant regions—a problem known as Navigation Among Movable Objects (NAMO). While various solutions to this problem have been developed, they often struggle to scale to highly cluttered environments. To address this, we propose NAMO-LLM, a sampling-based planner that searches over robot and obstacle configurations to compute feasible plans specifying which obstacles to move, where, and in what order. Its key novelty is a non-uniform sampling strategy guided by Large Language Models (LLMs) biasing the tree construction toward directions more likely to yield a solution. We show that NAMO-LLM is probabilistically complete and demonstrate through experiments that it efficiently scales to cluttered environments, outperforming related works in both runtime and plan quality.
I Introduction
Several planning algorithms have been developed to compute robot paths from an initial configuration to a desired one while avoiding obstacles [1, 2, 3, 4, 5]. However, in highly cluttered environments, obstacle-free paths to goal regions may not exist. In such cases, rather than reporting task failure, a more effective approach is to enable robots to interact with their environment by repositioning movable obstacles. This requires planning algorithms capable of determining which obstacles to move, in what order, and where to reposition them to create a feasible path to the goal; see Fig. 1. This planning problem, known as Navigation Among Movable Objects (NAMO) [6], has been proven to be NP-hard [7].
Several algorithms have been proposed to address NAMO problems. Early approaches relied on heuristic solutions which lacked completeness guarantees [8]. To overcome this, search-based planning methods have been introduced. These methods either provide completeness guarantees for specific subclasses of NAMO problems or sacrifice computational efficiency to ensure completeness across a broader range of scenarios [6, 9, 10, 11]. For instance, [6] addressed NAMO problems where disconnected free-space components can be connected by moving a single obstacle at a time. This formulation allows the problem to be represented as a graph structure, enabling the use of resolution-complete planners. This work was later extended to handle cases where at most $k$ obstacles must be moved, with each obstacle requiring only a single displacement [10]. A more general approach is proposed in [11] to solve non-monotone NAMO problems, where multiple obstacles may need to be repositioned—potentially more than once—to establish connectivity between disconnected free-space components. While this algorithm is complete under mild assumptions, its computational cost scales exponentially with the number of movable objects. Extensions of these approaches to account for physical interactions between objects are presented in [12, 13], while extensions to unknown environments are considered in [14, 15, 16, 17, 18, 19]. In addition to search-based methods, sampling-based planners have been developed to improve efficiency while still maintaining (probabilistic) completeness guarantees for general NAMO problems [20, 21, 22, 23]. The key idea in these approaches is to incrementally construct trees by sampling and adding nodes, where each node represents the environment’s configuration, and edges model feasible manipulation actions. However, their efficiency declines as the number of movable objects or the number of required obstacle relocations grows, due to their uniform sampling strategy.
Figure 1: A ground robot is tasked with reaching a destination behind the building block; see (1). All direct pathways are initially blocked (see Fig. 2). The robot autonomously removes obstacles to clear a path; see (2)–(4).
This paper addresses this limitation by designing a new sampling-based NAMO algorithm that efficiently scales to environments characterized by large numbers of movable obstacles and/or the need for multiple obstacle relocations to reach desired regions. Inspired by [22], our algorithm incrementally builds trees that explore the space of obstacle and robot configurations. Each tree node represents an environment configuration, while edges among nodes correspond to manipulation actions that relocate movable obstacles. At each iteration, a tree node and a manipulation action are sampled. Application of that action to the sampled tree node results in new tree node that is added to the tree if feasible. The key novelty of our work lies in a non-uniform sampling strategy that leverages Large Language Models (LLMs) to accelerate plan generation. Specifically, our sampling strategy directs the tree to grow towards LLM-recommended directions at certain times while allowing random exploration otherwise. We refer to the resulting planner as NAMO-LLM, for Navigation Among Movable Obstacles with Large Language Model guidance. We show that NAMO-LLM is probabilistically complete meaning that the probability of computing a feasible plan converges to one as the number of iterations goes to infinity. We conduct extensive comparative experiments against both search-based planners [18] and sampling-based planners relying on uniform sampling strategies [22], demonstrating that our method significantly reduces both the time required to compute the first feasible plan and the corresponding number of obstacles that need to be relocated. We also demonstrate NAMO-LLM on hardware experiments; see Fig. 1.
Related Works: In addition to the NAMO algorithms discussed earlier, our work is also related to research on rearrangement planning (RP) problems where robots manipulate movable obstacles [24, 25, 26, 27]. RP differs from NAMO because it considers predefined goal positions for the movable obstacles and typically does not specify a final goal for the robot. While our work primarily focuses on NAMO problems, we believe that our non-uniform sampling strategy can also be integrated with various sampling-based approaches for planning among movable obstacles, including RP and manipulation planning [28, 29]. However, a formal integration with these approaches is beyond the scope of this paper. Furthermore, several works design non-uniform sampling strategies to accelerate sampling-based planners [30, 31, 32, 33, 34, 35, 36]. For example, [34] learns a non-uniform sampling distribution from demonstrations. However, we emphasize that these works target different planning problems from ours, focusing on designing obstacle-free paths between an initial and a desired robot configuration (while also assuming that such paths exist). As such, they cannot be directly applied to our setting. It is worth noting that, unlike [32, 33, 34, 35, 36], our work does not require training neural networks; instead, ours leverages pre-trained LLMs.
Summary of Contributions: First, we introduce NAMO-LLM, a novel sampling-based planner designed to efficiently solve NAMO problems in highly cluttered environments. Second, to the best of our knowledge, this is the first integration of LLMs within a sampling-based NAMO planner, significantly accelerating plan generation. Third, we conduct extensive experiments demonstrating that our approach outperforms existing planning algorithms in terms of efficiency and solution quality. Fourth, we release an open-source implementation of our proposed algorithm [37].
II Problem Formulation
Robot Modeling: We consider a robot with the following kinematics: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ where $\mathbf{p}(t)\in{\mathcal{P}}$ and $\mathbf{u}(t)\in{\mathcal{U}}$ denote the robot configuration (e.g., position and orientation) and the control input at time step $t$, respectively. We assume that ${\mathbf{p}}(t)$ is known at all times and that the robot is holonomic and, therefore, it can move along any desired path. The robot is also equipped with a gripper, enabling it to manipulate one movable object at a time.
Figure 2: Top-view representation of the urban environment shown in Fig. 1. The set ${\mathcal{P}}$ is defined by the position of the center of the disk-shaped robot. Given ${\mathcal{N}}(0)$, the set ${\mathcal{O}}_{M}(0)$ includes only obstacles 1-6.
Environment Modeling: The robot operates in an environment $\Omega$ containing a set of known obstacles ${\mathcal{O}}=\{o_{1},\dots,o_{O}\}$. The obstacles are classified as movable obstacles, collected in a set ${\mathcal{O}}_{\text{movable}}\subseteq{\mathcal{O}}$, and non-movable ones. A movable object can be repositioned by the robot as long as it first grasps it. The configuration (e.g., position and orientation) of an obstacle $o_{i}\in{\mathcal{O}}$ at time $t$ is denoted by $c_{i}(t)\in{\mathcal{C}}_{i}$ where ${\mathcal{C}}_{i}$ denotes its configuration space. We also denote by $co_{i}(t)$ the region in the configuration space $\mathcal{P}$ of the robot that corresponds to configurations where the robot is in collision with obstacle $o_{i}$; see Fig. 2. Thus, the robot’s free space at time $t$, denoted by ${\mathcal{P}}_{\text{free}}(t)$, is defined as ${\mathcal{P}}_{\text{free}}(t)={\mathcal{P}}\setminus\mathcal{CO}(t)$ where $\mathcal{CO}(t)=\cup_{i=1}^{n_{O}}\{co_{i}(t)\}$. Notice that ${\mathcal{P}}_{\text{free}}(t)$ may consist of multiple connected components whose boundaries are determined by the configuration of the obstacles. We denote by ${\mathcal{N}}(t)\subseteq{\mathcal{P}}_{\text{free}}(t)$, the free-space component that ${\mathbf{p}}(t)$ lies in; see Fig. 2.
Robot Task: The robot is tasked with reaching a desired configuration starting from an initial one ${\mathbf{p}}(0)$ while avoiding collisions with obstacles. Hereafter, we denote by ${\mathcal{P}}_{g}\subseteq{\mathcal{P}}$ a set collecting all desired robot configurations. We consider cases where all configurations in ${\mathcal{P}}_{g}$ are initially unreachable due to the initial configuration of obstacles, i.e., ${\mathcal{P}}_{g}\cap{\mathcal{N}}(0)=\emptyset$. In such cases, the robot must determine which movable obstacles need to be relocated and where they should be placed. To formalize this, we define the set of manipulable obstacles at time $t$ as $\mathcal{O}_{M}(t)\subseteq{\mathcal{O}}_{\text{movable}}$ which includes all movable obstacles $o_{i}$ that the robot can directly reach, i.e., $\mathcal{O}_{M}(t)=\{o_{i}\in{\mathcal{O}}_{\text{movable}}\mid co_{i}(t)\cap{\mathcal{N}}(t)\neq\emptyset\}$; see also Fig. 2. Next, we define manipulation actions of the form $a(c_{i}(t),c_{i}^{\prime})$ which involve moving a manipulable object $o_{i}\in\mathcal{O}_{M}(t)$, from its current configuration $c_{i}(t)$ to a new configuration $c_{i}(t^{\prime})=c_{i}^{\prime}$, for some $t^{\prime}>t$. In other words, the application of this action requires the robot to approach the obstacle’s current location (determined by $c_{i}(t)$), grasp it, and relocate it to the new configuration $c_{i}^{\prime}$. Thus, our goal is to compute a plan $\tau$, i.e., a sequence of manipulation actions so that eventually ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t)\neq\emptyset$. We defined this plan as $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, where $\tau(k)$ refers to the $k$-th manipulation action, $k\in\{0,1,\dots,H\}$, for some horizon $H\geq 0$. Note that the index $k$ is different from the time step $t$, as the application of an action $\tau(k)$ may require multiple time steps $t$. The problem that this paper addresses can be summarized as follows:
Problem 1
Given an initial robot configuration ${\mathbf{p}}(0)$, a set ${\mathcal{P}}_{g}$ of desired configurations, and initial obstacle configurations $c_{i}(0)$, compute a manipulation plan $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, so that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$.
III Proposed Planning Algorithm
In this section, we present NAMO-LLM, our proposed planner to address Problem 1. In Section III-A, we present a sampling-based planner to design plans $\tau$. In Section III-B, we present a non-uniform sampling strategy that accelerates the plan synthesis phase by leveraging Large Language Models (LLM). In Section III-C, we discuss the input prompt to the LLM guiding the sampling strategy. In Section III-D, we show that NAMO-LLM is probabilistically complete.
Input: (i) initial robot configuration ${\mathbf{p}}(0)$; (ii) initial obstacle configuration ${\mathbf{c}}(0)$; (iii) Set ${\mathcal{O}}_{\text{movable}}$ ; (iv) goal set ${\mathcal{X}}_{g}$
Output: The manipulation plan $\tau$
1 Compute ${\mathcal{N}}(0)$;
2 ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ ;
3 Initialize ${\mathcal{V}}=\{{\mathbf{q}}_{0}\},~{}{\mathcal{E}}=\emptyset,~{}{\mathcal{X}}_{g}=\emptyset,~{}m=1$ ;
4 while $m<m_{\text{max}}$ do
5       Sample a node ${\mathbf{q}}_{\text{rand}}\in{\mathcal{V}}$ using $f_{\mathcal{V}}$;
6       $[{\mathbf{q}}_{\text{new}},a(c_{i}(t),c_{i}^{\prime})]=\texttt{TreeExpansion}({\mathbf{q}}_{\text{rand}})$;
7       if  ${\mathbf{q}}_{\text{new}}\neq\varnothing$ then
8             ${\mathcal{V}}\leftarrow{\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$, ${\mathcal{E}}\leftarrow{\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$;
9             $A(({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}}))=a(c_{i}(t),c_{i}^{\prime})$;
10             Extract ${\mathcal{N}}(t^{\prime})$ from ${\mathbf{q}}_{\text{new}}=[\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$;
11             if  ${\mathcal{P}}_{g}\cap\mathcal{N}(t^{\prime})\neq\emptyset$  then
12                   ${\mathcal{X}}_{g}\leftarrow{\mathcal{X}}_{g}\cup\{{\mathbf{q}}_{\text{new}}\}$;
13
14      Update iteration index: $m=m+1$;
15
16if ${\mathcal{X}}_{g}\neq\emptyset$ then
17       Pick a node ${\mathbf{q}}_{g}\in{\mathcal{X}}_{g}$;
18       Compute path $\tau_{{\mathbf{q}}}$ over ${\mathcal{G}}$ from ${\mathbf{q}}(0)$ to ${\mathbf{q}}_{g}$;
19       Using $A$ and $\tau_{{\mathbf{q}}}$, compute the manipulation plan $\tau$ ;
20
21else
22       No plan $\tau$ has been found;
23
Algorithm 1 NAMO-LLM
1 Compute the set of manipulable obstacles ${\mathcal{O}}_{M,\text{rand}}(t)$ ;
2 Sample a manipulable obstacle $o_{i}$ from a given distribution $f_{{\mathcal{O}}_{M}}$ ;
3 Sample a random configuration $c_{i}^{\prime}$ from $f_{{\mathcal{C}}_{i}}$;
4 if $co_{i}^{\prime}\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ and $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$  then
5       if $c_{i}(t)\to c^{\prime}_{i}$ is collision-free  then
6             Initialize ${\mathbf{c}}(t^{\prime})={\mathbf{c}}(t)$ and update $c_{i}(t^{\prime})=c_{i}^{\prime}$ in ${\mathbf{c}}(t^{\prime})$;
7             $\mathcal{P}_{\text{free}}(t^{\prime})=\mathcal{P}_{\text{free}}(t)\cup co_{i}(t)\setminus co_{i}^{\prime}$ ;
8             Given ${\mathbf{c}}(t^{\prime})$, select $\mathcal{N}(t^{\prime})$ ;
9             ${\mathbf{q}}_{\text{new}}=(\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime}))$ ;
10             return ${\mathbf{q}}_{\text{new}}$, $a(c_{i}(t),c_{i}^{\prime})$
11      else
12             return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
13else
14       return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
Algorithm 2 TreeExpansion(${\mathbf{q}}_{\text{rand}}$)
III-A Sampling-based Planning
We propose a sampling-based planner to address Problem 1 summarized in Alg. 1 The proposed algorithm incrementally builds a tree that explores the configuration space of the movable objects and the robot’s free space components to compute a manipulation plan $\tau$ as defined in Section II. We denote the constructed tree by ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where ${\mathcal{V}}$ and ${\mathcal{E}}$ is a set of nodes and edges, respectively. The set ${\mathcal{V}}$ contains states of the form ${\mathbf{q}}(t)=[\mathcal{N}(t),{\mathbf{c}}(t)]$, where ${\mathbf{c}}(t)$ is a vector stacking the configuration of all obstacles, i.e., ${\mathbf{c}}(t)=[c_{1}(t),c_{2}(t),\dots,c_{n}(t)]$ and ${\mathcal{N}}(t)$ denotes the free-space component that ${\mathbf{p}}(t)$ lies in. The root of the tree models the initial configuration of the obstacles and the robot, i.e., ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ [lines 1-1, Alg. 1]. The tree is initialized so that ${\mathcal{V}}=\{{\mathbf{q}}(0)\}$ and ${\mathcal{E}}=\emptyset$ [line 1, Alg. 1]. Also, the tree is built incrementally by sampling new states ${\mathbf{q}}_{\text{new}}$ that are added to the tree structure [line 1, Alg. 1]. The algorithm terminates after a user-specified number of iterations $m_{\text{max}}$ and returns a feasible plan $\tau$ (if it has been found). In what follows, we present a detailed description of our method.
At every iteration $m$ of Alg. 1, we perform the following two steps. First, we sample from a given discrete mass function $f_{{\mathcal{V}}}:{\mathcal{V}}\rightarrow[0,1]$ a tree node denoted by ${\mathbf{q}}_{\text{rand}}(t)=[{\mathcal{N}}_{\text{rand}}(t),{\mathbf{c}}_{\text{rand}}(t)]\in{\mathcal{V}}$ [line 1, Alg. 1]. The mass function $f_{{\mathcal{V}}}$ defines the probability of selecting a node ${\mathbf{q}}\in{\mathcal{V}}$. Second, we sample a new node ${\mathbf{q}}_{\text{new}}$ that will be added to the tree; see Alg. 2. Specifically, given ${\mathbf{q}}_{\text{rand}}$, we compute the corresponding set of manipulable obstacles denoted by $\mathcal{O}_{M,\text{rand}}(t)$, [line 2, Alg. 2]. Then, we sample from a discrete mass function $f_{{\mathcal{O}}_{M}}:{\mathcal{O}}_{M,\text{rand}}\rightarrow[0,1]$ a manipulable obstacle $o_{i}\in\mathcal{O}_{M}(t)$ [line 2, Alg. 2]. The mass function $f_{{\mathcal{O}}_{M}}$ defines the probability of selecting a manipulable obstacle $o_{i}\in{\mathcal{O}}_{M,\text{rand}}$. Given the selected manipulable obstacle $o_{i}$, we sample a new configuration $c_{i}^{\prime}\in{\mathcal{C}}_{i}$ using a density function $f_{\mathcal{C}_{i}}:\mathcal{C}_{i}\rightarrow[0,\infty)$ [line 2, Alg. 2]. Designing the functions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ is deferred to Section III-B.
Figure 3: Illustration of the free-space components ${\mathcal{N}}(t)$ (left column) and ${\mathcal{N}}^{1}(t)$ (right column) in two environments.
Next, we check if a collision-free path exists for the robot to move $o_{i}$ from $c_{i}(t)$ to $c_{i}^{\prime}$ [line 2, Alg. 2]. Two necessary conditions must be met for $c_{i}^{\prime}$ to admit such a collision-free path; if they are not satisfied then such a path does not exist [line 2, Alg. 2]. First, $c_{i}^{\prime}$ must not intersect other obstacles, i.e., $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}^{\text{rand}}(t)\setminus co_{i}(t)$, where ${\mathcal{P}}_{\text{free}}^{\text{rand}}(t)$ is computed using ${\mathbf{q}}_{\text{rand}}$. Second, $co_{i}^{\prime}$ must not lie in a component of ${\mathcal{P}}_{\text{free}}$ that becomes accessible only if an obstacle $o_{j}\neq o_{i}$ is relocated first. To formalize this, we define ${\mathcal{P}}^{i}_{\text{free}}(t)$ as the robot’s free space if obstacle $o_{i}$ is removed from the environment, i.e., ${\mathcal{P}}^{i}_{\text{free}}(t)={\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$. Correspondingly, ${\mathcal{N}}^{i}(t)$ denotes the component of ${\mathcal{P}}^{i}_{\text{free}}(t)$ containing the robot’s position ${\mathbf{p}}(t)$; see Fig. 3. Thus, the second condition requires $co_{i}^{\prime}$ to belong to ${\mathcal{N}}^{i}_{\text{rand}}(t)$ (i.e., to ${\mathcal{N}}^{i}(t)$ computed using ${\mathbf{q}}_{\text{rand}}$).
If the above two conditions hold, then there may exist such a collision-free path for $o_{i}$. This path is defined as a sequence of obstacle configurations, i.e., $\boldsymbol{\zeta}=c_{i}(t),c_{i}(t+1),\dots,c_{i}(t+s),\dots,c_{i}(t^{\prime})$ such that (i) $c_{i}(t^{\prime})=c_{i}^{\prime}$ and (ii) $co_{i}(t+s)\in{\mathcal{N}}^{i}(t)$, for all $s\in\{1,2,\dots,t^{\prime}-t\}$. This can be checked using any existing motion planner such as RRT or A* [3]. If such a path $\boldsymbol{\zeta}$ exists, this means the obstacle $o_{i}$ can be moved to $c_{i}^{\prime}$ giving rise to ${\mathbf{c}}(t^{\prime})$ that differs from ${\mathbf{c}}(t)$ only in the configuration of the obstacle $o_{i}$ [line 2, Alg. 2]. In general, ${\mathcal{N}}^{i}(t)$ may split into multiple components once $o_{i}$ is moved to $c_{i}^{\prime}$, each adjacent to $co_{i}^{\prime}$. The robot can be in any of these components, and we randomly select one to be ${\mathcal{N}}(t^{\prime})$ [line 2, Alg. 2].
Given the selected free-space component ${\mathcal{N}}(t^{\prime})$ and ${\mathbf{c}}(t^{\prime})$, we construct the node ${\mathbf{q}}_{\text{new}}=[{\mathcal{N}}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$ [line 2, Alg. 2] and the sets of nodes and edges are updated as ${\mathcal{V}}={\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$ and ${\mathcal{E}}={\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$, respectively [lines 1, Alg. 1]. Also, once an edge $({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ is added to the tree, we store the corresponding manipulation $a(c_{i}(t),c_{i}^{\prime})$ that can enable it. We formalize this by iteratively constructing a function $A$ which returns the corresponding action for each edge in ${\mathcal{E}}$ [line 1, Alg. 1]. If a collision-free path $\boldsymbol{\zeta}$ does not exist or if $c_{i}^{\prime}$ is not a valid configuration, then we set $q_{\text{new}}=\varnothing$ and the iteration $m$ will terminate without adding any new nodes to the tree [lines 2-2, Alg. 2].
Once a node ${\mathbf{q}}_{\text{new}}$ is added to the tree, we check if ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t^{\prime})\neq\emptyset$. If that holds, it means that the robot can reach its goal configuration set ${\mathcal{P}}_{g}$. The tree nodes that satisfy this property are collected in a set ${\mathcal{X}}_{g}\subseteq{\mathcal{V}}$ [line 1, Alg. 1] that is initialized as ${\mathcal{X}}_{g}=\emptyset$ [line 1, Alg. 1]. Once the algorithm terminates, we select a tree node ${\mathbf{q}}_{g}\in{\mathcal{X}}_{g}$ (if it is non-empty) [lines 1-1, Alg. 1]. Then, we compute the path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(H+1)$ over ${\mathcal{G}}$ connecting ${\mathbf{q}}_{g}$ to the tree root ${\mathbf{q}}(0)$, i.e., $\tau_{{\mathbf{q}}}(0)={\mathbf{q}}(0)$, $\tau_{{\mathbf{q}}}(H+1)={\mathbf{q}}_{g}$, and $\tau_{{\mathbf{q}}}(k)\in{\mathcal{V}}$, $\forall k\in\{1,\dots,H+1\}$ [line 1, Alg. 1]. Using $\tau_{{\mathbf{q}}}$ and the function $A$, we can extract the plan $\tau=\tau(0),\dots,\tau(H)$ modeling the sequence of manipulation actions that need to be applied to solve Problem 1 [line 1, Alg. 1]. The node ${\mathbf{q}}_{g}$ can be selected randomly or based on any user-defined criterion such as the horizon $H$ of $\tau$ or the distance that needs to be traveled to execute $\tau$; see also Rem. III.1.
Remark III.1 (Execution of $\tau$)
Each action $\tau(k)=a(c_{i},c_{i}^{\prime})$ in $\tau$ involves moving an object $o_{i}$ from its current configuration $c_{i}$ to a new configuration $c_{i}^{\prime}$. Execution of this action should be in accordance with the corresponding path $\tau_{{\mathbf{q}}}$. This means that the robot should grasp $o_{i}$, while being in ${\mathcal{N}}$ as determined by $\tau_{{\mathbf{q}}}(k)=[{\mathcal{N}},{\mathbf{c}}]$, and drop-it off at $c_{i}^{\prime}$ while being in ${\mathcal{N}}^{\prime}$ determined by $\tau_{{\mathbf{q}}}(k+1)=[{\mathcal{N}}^{\prime},{\mathbf{c}}^{\prime}]$. According to line 2 in Alg. 2, there must exist a collision-free path for the robot to execute this movement. Also, observe that, as in [22], our planner decouples the computation of the robot’s motion from the computation of obstacle movements. Specifically, it does not determine the exact grasp poses for the objects of interest but instead tracks the free-space component of the robot.
Figure 4: Input prompt for GPT-4o and $K=1$ in Case Study V (Sec. IV).
III-B LLM-guided Sampling Strategy
The sampling-based planner above relies on the distributions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ to generate ${\mathbf{q}}_{\text{new}}$. While these can be uniform, as e.g., in [22], such strategies often lead to slow plan synthesis, particularly as environmental complexity (e.g., the number of movable objects) increases. To accelerate the planning process, we propose non-uniform mass functions $f_{{\mathcal{V}}}$ and $f_{{\mathcal{O}}_{M}}$ that bias sampling toward directions where a feasible solution is more likely to be found. Our approach uses LLMs to identify these promising directions.
First, we define $f_{{\mathcal{V}}}$ as follows:
$f_{{\mathcal{V}}}({\mathbf{q}}\mid{\mathcal{V}},{\mathcal{V}}_{v})=\begin{cases}\frac{p_{\text{rand}}}{|{\mathcal{V}}_{v}|}+\frac{1-p_{\text{rand}}}{|{\mathcal{V}}|},&\text{if}~{}{\mathbf{q}}\in{\mathcal{V}}_{v}\\ \frac{1-p_{\text{rand}}}{|{\mathcal{V}}|},&\text{otherwise}\end{cases}$ (1)
where $p_{\text{rand}}\in[0,1)$ is a user-specified parameter and ${\mathcal{V}}_{v}\subseteq{\mathcal{V}}$; ${\mathcal{V}}_{v}$ will be defined later once $f_{{\mathcal{O}}_{M}}$ is defined. According to (1), a node ${\mathbf{q}}\in{\mathcal{V}}_{v}$ is selected to be the node ${\mathbf{q}}_{\text{rand}}$ with probability $p_{\text{rand}}$ while any node ${\mathbf{q}}\in{\mathcal{V}}$ is selected with probability $1-p_{\text{rand}}$. Notice that if $p_{\text{rand}}=0$, then $f_{{\mathcal{V}}}$ models uniform sampling. As $p_{\text{rand}}$ increases, the sampling strategy becomes more biased toward selecting nodes ${\mathbf{q}}\in{\mathcal{V}}_{v}$.
Second, we define the function $f_{{\mathcal{O}}_{M}}$ as follows:
$f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})=\begin{cases}p_{\text{obs}}+\frac{1}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}=o_{\text{LLM}},\\ \frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}\in\mathcal{O}_{M,\text{rand}},\end{cases}$ (2)
where $p_{\text{obs}}\in[0,1)$ is a user-specified parameter and $o_{\text{LLM}}$ is manipulable obstacle, i.e., $o_{\text{LLM}}\in{\mathcal{O}}_{M}$, that is recommended by an LLM to relocate. A discussion on how to prompt an LLM to select $o_{\text{LLM}}$ is deferred to Section III-C. According to (2), the obstacle $o_{\text{LLM}}\in{\mathcal{O}}_{M}$ is selected for relocation with probability $p_{\text{obs}}$ while any obstacle $o_{i}\in{\mathcal{O}}_{M}$ is selected with probability $1-p_{\text{obs}}$. The probability of selecting $o_{i}\notin{\mathcal{O}}_{M}$ is 00. Observe that if $p_{\text{obs}}=0$, then $f_{{\mathcal{O}}_{M}}$ models uniform sampling. As $p_{\text{obs}}$ increases, the sampling strategy becomes more biased toward selecting $o_{\text{LLM}}$. Also, we note that it is possible that the LLM may fail to select a manipulable obstacle or it may generate nonsensical information. In such cases, we set $p_{\text{obs}}=0$ for that iteration so that a manipulable obstacle is selected randomly; see Section III-C.
Next, we define the set ${\mathcal{V}}_{v}$ used for the construction of $f_{{\mathcal{V}}}$ in (1). First, we define a function $v:{\mathcal{V}}\rightarrow\mathbb{N}$ assigning a positive score to each node ${\mathbf{q}}\in{\mathcal{V}}$. This function is initialized as $v({\mathbf{q}}(0))=0$. Second, we show how the score of a newly added node ${\mathbf{q}}_{\text{new}}$ is generated. If ${\mathbf{q}}_{\text{new}}$ was constructed by relocating $o_{\text{LLM}}$, then we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})+1$. Otherwise, we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})$. Informally, $v({\mathbf{q}}_{\text{new}})$ counts how many nodes along the path connecting ${\mathbf{q}}_{\text{new}}$ to ${\mathbf{q}}(0)$ have been constructed based on LLM recommendations. Third, we construct the set ${\mathcal{V}}_{v}$ so that it collects the nodes with the highest score $v$, i.e., ${\mathcal{V}}_{v}=\{{\mathbf{q}}\in{\mathcal{V}}~{}|~{}v({\mathbf{q}})=\max_{{\mathbf{q}}^{\prime}\in{\mathcal{V}}}v({\mathbf{q}}^{\prime})\}$.
Finally, we select a uniform density function for $f_{{\mathcal{C}}_{i}}$. Non-uniform density functions $f_{{\mathcal{C}}_{i}}$ can also be implemented similarly, e.g., by querying an LLM to select configurations $c_{i}^{\prime}$ for which collision-free paths exist; for simplicity, we abstain from this design choice.
III-C Prompt Structure
In this section, we describe how we query a pre-trained LLM to select a manipulable obstacle $o_{\text{LLM}}$. The input prompt to the LLM consists of three parts (see Fig.4): (1) Task description which provides information about the environment’s structure (i.e., the location of all movable and non-movable obstacles and the goal region) and specifies that the robot’s task is to remove obstacles so that the goal region is reachable; (2) Rules that the robot must follow to accomplish its task. Specifically, these rules state that the robot can only remove manipulable obstacles, should minimize the number of obstacles removed and should avoid picking manipulable obstacles that are blocked by other obstacles (see e.g., Case Study VI in Section IV); (3) Output format which defines how the LLM’s response should be formatted for compatibility with Alg. 2. In particular, the LLM is required to generate $K\geq 1$ obstacles in the form $\text{obs}_{i}$ where $i$ is a unique identifier of each obstacle. Among the generated obstacles, one is selected randomly and denoted by $o_{\text{LLM}}$ in (2). Note that the LLM may make content and format errors. A content error occurs when the selected obstacle is not manipulable, and a format error occurs when the LLM’s output does not comply with the required format. If either of these errors occurs, we set $p_{\text{obs}}=0$.
III-D Probabilistic Completeness
In what follows, we show that NAMO-LLM is probabilistically complete. To formally state this result, we first introduce the notion of path clearance [22]. Consider any path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),\dots,\tau_{{\mathbf{q}}}(k),\dots,\tau_{{\mathbf{q}}}(H+1)$ computed by Alg. 1 yielding a plan $\tau$ solving Problem 1. We say that $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$ if any alternative path $\tilde{\tau}_{{\mathbf{q}}}=\tilde{\tau}_{{\mathbf{q}}}(0),\dots,\tilde{\tau}_{{\mathbf{q}}}(k),\dots,\tilde{\tau}_{{\mathbf{q}}}(H+1)$ satisfying the following conditions (i)-(ii) for all $k\in\{1,\dots,H\}$ also corresponds to a feasible plan $\tilde{\tau}$. (a) The transitions from $\tau_{{\mathbf{q}}}(k-1)$ to $\tau_{{\mathbf{q}}}(k)$ and from $\tilde{\tau}_{{\mathbf{q}}}(k-1)$ to $\tilde{\tau}_{{\mathbf{q}}}(k)$ must involve relocating the same manipulable obstacle $o_{i}$, albeit potentially to different configurations. (b) Let $c_{i}^{k}$ and $\tilde{c}_{i}^{k}$ denote the configurations of that obstacle $o_{i}$ in $\tau_{{\mathbf{q}}}(k)$ and $\tilde{\tau}_{{\mathbf{q}}}(k)$, respectively. These configurations must satisfy $\lVert c_{i}^{k}-\tilde{c}_{i}^{k}\rVert\leq\epsilon$ for all $k$, for some $\epsilon>0$. In what follows, we refer to the satisfaction of conditions (i)–(ii) concisely as $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$.
Theorem III.2
Assume that there exists a feasible path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),...,\tau_{{\mathbf{q}}}(H+1)$ with clearance $\epsilon>0$. If $p_{\text{rand}}\in[0,1)$, $p_{\text{obs}}\in[0,1)$, and the motion planner used to check existence of collision-free paths between obstacle configurations in line 2 of Alg. 2 is complete, then the probability that NAMO-LLM will find a feasible path approaches $1$ as $m$ goes to $\infty$.
Proof:
To show this result, we need first to introduce the following notations and definitions. Let ${\mathcal{G}}_{m}=\{{\mathcal{V}}_{m},{\mathcal{E}}_{m}\}$ denote the tree constructed by Algorithm 1 at the end of iteration $m$. Since the plan $\tau_{{\mathbf{q}}}$ has clearance $\epsilon>0$, there exist alternative feasible plans $\tilde{\tau}_{{\mathbf{q}}}$ such that $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$ for all $k$. Throughout the proof, any reference to a state $\tilde{\tau}_{{\mathbf{q}}}(k)$ assumes this $\epsilon$-closeness to the plan $\tau_{{\mathbf{q}}}$.
We also define the following two random variables that will be used to model the sampling process and the computation of a feasible plan. First, we define Bernoulli random variables $X_{m}$ that are equal to $1$ if the following holds for any state $\tilde{\tau}_{{\mathbf{q}}}(k)$ and $k\in\{0,1,\dots,H+1\}$ at iteration $m$: (i) a state $\tilde{\tau}_{{\mathbf{q}}}(k)$ is added to the tree (i.e., $\tilde{\tau}_{{\mathbf{q}}}(k)\in{\mathcal{V}}_{m}$) via an edge from a predecessor $\tilde{\tau}_{{\mathbf{q}}}(k-1)\in{\mathcal{V}}_{m^{\prime}}$ for some $m^{\prime}<m$ (i.e., $(\tilde{\tau}_{{\mathbf{q}}}(k-1),\tilde{\tau}_{{\mathbf{q}}}(k))\in{\mathcal{E}}_{m}$), and (ii) no state $\tilde{\tau}_{{\mathbf{q}}}^{\prime}(k)$ satisfying $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}^{\prime}(k)\rVert\leq\epsilon$ was already present in the tree prior to iteration $m$ (i.e., $\tilde{\tau}_{{\mathbf{q}}}^{\prime}(k)\notin{\mathcal{V}}_{m^{\prime}}$ for all $m^{\prime}<m$). Additionally, once a state $\tilde{\tau}_{{\mathbf{q}}}(H+1)$ satisfying both (i) and (ii) is added at an iteration $m$, no future iterations can satisfy condition (ii) for any $k$. In this case, by convention, we set $X_{m^{\prime}}=1$, for all $m^{\prime}>m$ (i.e., the probability of success of $X_{m^{\prime}}$ is $1$).
Second, we define the random variable $Y_{m}$ counting the number of successes of the Bernoulli variables $X_{1},X_{2},\dots,X_{m}$, i.e., $Y_{m}=\sum_{i=1}^{m}X_{i}$. By construction of $X_{i}$, if $Y_{m}>H$, then a state $\tilde{\tau}_{{\mathbf{q}}}(H+1)$ has been added to ${\mathcal{V}}_{m}$ indicating that a solution has been found. Therefore, to show probabilistic completeness, it suffices to show that the probability $\mathbb{P}(Y_{m}\geq H+1)$ approaches $1$ as $m\to\infty$.
Using the Chebyshev inequality, we can bound the “failure” probability, i.e., the probability that fewer than $H$ successful expansions occurred in $m$ iterations (which means the goal has not been reached), as follows:
$\mathbb{P}(Y_{m}\leq H)$ $\leq\mathbb{P}\Bigl{(}|Y_{m}-\mathrm{E}(Y_{m})|\geq\mathrm{E}(Y_{m})-H\Bigl{)}$
$\leq\frac{\mathrm{Var}(Y_{m})}{(\mathrm{E}(Y_{m})-H)^{2}}$ (3)
The expectation $\mathrm{E}(Y_{m})$ and variance $\mathrm{Var}(Y_{m})$ of $Y_{m}$ are defined as follows
$\mathrm{E}(Y_{m})=\sum_{i=1}^{m}\mathrm{E}(X_{i})=\sum_{i=1}^{m}p_{i}$ (4)
and
$\mathrm{Var}(Y_{m})=\sum_{i=1}^{m}\mathrm{Var}(X_{i})+2\sum_{1\leq i<j\leq m}\mathrm{Cov}(X_{i},X_{j})$
$=\sum_{i=1}^{m}p_{i}(1-p_{i})+2\sum_{1\leq i<j\leq m}\Bigl{[}\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\Bigr{]},$ (5)
where $p_{i}$ is the probability of success of $X_{i}$.
In what follows, we show (a) $\lim_{m\to\infty}E(Y_{m})=\infty$ and (b) $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$. Using (b), we can re-write (III-D) as follows:
$\mathbb{P}(Y_{m}\leq H)\leq\frac{\mathrm{Var}(Y_{m})}{(\mathrm{E}(Y_{m})-H)^{2}}\leq\frac{3\mathrm{E}(Y_{m})}{(\mathrm{E}(Y_{m})-H)^{2}}.$ (6)
Then combining (a) with (6), we get that $\lim_{m\to\infty}\mathbb{P}(Y_{m}\leq H)=0$. This equivalently means $\lim_{m\to\infty}\mathbb{P}(Y_{m}\geq H+1)=1$ completing the proof.
Proof of (a): Here we show that $\lim_{m\to\infty}E(Y_{m})=\infty$. Step 1: To show this, we first show the probability of success of $X_{m}$, denoted by $p_{m}$, is strictly positive, i.e., $p_{m}>0$, for all $m$. This result will later be used to establish a lower bound on $\mathbb{P}(Y_{m}\geq H+1)$. Showing this consists of the following two parts.
Part 1: First, let $\hat{{\mathcal{V}}}_{m}\subseteq{\mathcal{V}}_{m}$ collect all states $\tilde{\tau}_{{\mathbf{q}}}(k)\in{\mathcal{V}}_{m}$ for which there does not exist a state $\tilde{\tau}_{{\mathbf{q}}}(k^{\prime})$ with $k^{\prime}>k$ that is already in the set ${\mathcal{V}}_{m}$. A necessary requirement for $X_{m}=1$ is to select a state from $\hat{{\mathcal{V}}}_{m-1}$ to be ${\mathbf{q}}_{\text{rand}}$. By construction of $f_{{\mathcal{V}}}({\mathbf{q}}\mid{\mathcal{V}},{\mathcal{V}}_{v})$, the probability $p_{1}^{m}=f_{{\mathcal{V}}}({\mathbf{q}}\mid{\mathcal{V}},{\mathcal{V}}_{v})$ of selecting a state ${\mathbf{q}}_{\text{rand}}\in\hat{{\mathcal{V}}}_{m-1}$ is strictly positive. Specifically, we have that $p_{1}^{m}\geq\min\{\frac{p_{\text{rand}}}{|{\mathcal{V}}_{v}|}+\frac{1-p_{\text{rand}}}{|{\mathcal{V}}_{m-1}|},\frac{1-p_{\text{rand}}}{|{\mathcal{V}}_{m-1}|}\}=\frac{1-p_{\text{rand}}}{|{\mathcal{V}}_{m-1}|}\geq\frac{1-p_{\text{rand}}}{m}$ where the last inequality is due to $|{\mathcal{V}}_{m-1}|\leq{m}$. Observe that $p_{2}^{m}>0$ since by assumption $p_{\text{rand}}<1$.
Part 2: Second, assume that a node $\tilde{\tau}_{{\mathbf{q}}}(k)$ from $\hat{{\mathcal{V}}}_{m}$ has been selected to be ${\mathbf{q}}_{\text{rand}}$. Under this assumption, there exist nodes ${\mathbf{q}}_{\text{new}}$ that can be added to the tree satisfying conditions (i)-(ii) required for $X_{m}=1$. Specifically, all nodes $\tilde{\tau}_{{\mathbf{q}}}(k+1)$ are such candidate nodes. Next, we show that the probability $p_{2}^{m}$ of sampling a state $\tilde{\tau}_{{\mathbf{q}}}(k+1)$ to be $q_{\text{new}}$ is strictly positive, i.e., $p_{2}^{m}>0$, for all $m$. Given ${\mathbf{q}}_{\text{rand}}=\tilde{\tau}_{{\mathbf{q}}}(k)$, sampling a state ${\mathbf{q}}_{\text{new}}=\tilde{\tau}_{{\mathbf{q}}}(k+1)$ will occur under the following two requirements (as per the $\epsilon$-clearance definition). First, at state ${\mathbf{q}}_{\text{rand}}=\tilde{\tau}_{{\mathbf{q}}}(k)$ the obstacle $o_{i}$ that will be selected to relocate is the one selected in the transition from $\tau_{{\mathbf{q}}}(k)$ to $\tau_{{\mathbf{q}}}(k+1)$. By construction of $f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})$, since the obstacle $o_{i}$ is manipulable (by construction of $\tau_{{\mathbf{q}}}$) this will occur with probability $p_{o_{i}}^{m}\geq\min\{p_{\text{obs}}+\frac{1}{|\mathcal{O}_{M,\text{rand}}|},\frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|}\}$. Observe that since $p_{\text{obs}}<1$, we have that $p_{o_{i}}^{m}>0$ for all $m$. Second, assuming that $o_{i}$ will be selected for relocation, its new configuration, denoted by $\tilde{c}_{i}^{k+1}$, should satisfy $\lVert\tilde{c}_{i}^{k+1}-c_{i}^{k+1}\rVert{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\leq\epsilon}$ where $c_{i}^{k+1}$ is the configuration of $o_{i}$ at state $\tau_{{\mathbf{q}}}(k+1)$. Since by assumption $\epsilon>0$ and $f_{{\mathcal{C}}_{i}}$ is a uniform density function, we get that the probability $p_{c_{i}}^{m}$ of sampling a $\tilde{c}_{i}^{k+1}$ that satisfies this requirement is also strictly positive for all $m$. Thus, we get that $p_{2}^{m}=p_{o_{i}}^{m}p_{c_{i}}^{m}>0$. Consequently, there must exist a positive constant $p_{2}^{\text{min}}$ such that $p_{2}^{m}\geq p_{2}^{\text{min}}$ for all $m\geq 0$ We note also since a complete planner is used to check if there exists a collision-free path reaching $\tilde{c}_{i}^{k+1}$ from $\tilde{c}_{i}^{k}$ (i.e., the configuration of the obstacle $o_{i}$ in $\tilde{\tau}_{{\mathbf{q}}}(k)$), then the sampled configuration $\tilde{c}_{i}^{k+1}$ will be never be rejected in line 2 incorrectly (i.e., if such a collision-free path exists).
Combining Part 1 and 2, we get that the probability $p_{m}$ of success of $X_{m}$ is strictly positive since $p_{m}=p_{1}^{m}p_{2}^{m}\geq\frac{(1-p_{\text{rand}})p_{2}^{m}}{m}\geq\frac{(1-p_{\text{rand}})p_{2}^{\text{min}}}{m}>0.$
Step 2: Given the result from Step 1, we can re-write the expected value of $Y_{m}$ in (4) as:
$E(Y_{m})=\sum_{i=1}^{m}p_{1}^{i}p_{2}^{i}\geq\sum_{i=1}^{m}\frac{(1-p_{\text{rand}})p_{2}^{\text{min}}}{i}=\alpha\sum_{i=1}^{m}\frac{1}{i}$
where $\alpha=(1-p_{\text{rand}})p_{2}^{\text{min}}>0$ is a positive constant across all iterations $m$. Then,
$\mathrm{E}(Y_{m})\geq\alpha\sum_{i=1}^{m}\frac{1}{i}$
When $m\to\infty$, $\sum_{i=1}^{m}\frac{1}{i}\to\infty$. Therefore, we get $\mathrm{E}(Y_{m})\to\infty$ completing the proof of part (a).
Proof of (b): Here we show that $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$. Recall from (III-D) that the variance of $Y_{m}$ is given by:
$\mathrm{Var}(Y_{m})$ $=\sum_{i=1}^{m}p_{i}(1-p_{i})+2\sum_{1\leq i<j\leq m}\Bigl{[}E(X_{i}X_{j})-p_{i}p_{j}\Bigr{]}$ (7)
First, we focus on the term $\sum_{i=1}^{m}p_{i}(1-p_{i})$ in (7). Since $0<p_{i}\leq 1$, we have that
$\sum_{i=1}^{m}p_{i}(1-p_{i})<\sum_{i=1}^{m}p_{i}=\mathrm{E}(Y_{m})$
Second, we focus on the term $\sum_{1\leq i<j\leq m}\Bigl{[}E(X_{i}X_{j})-p_{i}p_{j}\Bigr{]}$ in (7). Since $X_{i}X_{j}\leq X_{i}$, we have that:
$\mathrm{E}(X_{i}X_{j})\leq\mathrm{E}(X_{i})=p_{i}$
Hence,
$\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\leq p_{i}(1-p_{j})<p_{i}$
Then,
$2\sum_{1\leq i<j\leq m}\Bigl{[}\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\Bigr{]}<2\sum_{i=1}^{m}p_{i}=2\mathrm{E}(Y_{k})$
Thus we get $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$ completing the proof of part (b). ∎
IV Experiments
In this section, we present extensive numerical experiments to evaluate the efficiency of our method. Specifically, we demonstrate that our approach can effectively solve NAMO problems involving hundreds of movable obstacles and requiring the manipulation of multiple obstacles, significantly outperforming recent search-based methods [18] and sampling-based planners relying on uniform sampling [22] in terms of both runtime and plan quality. Hardware demonstrations of NAMO-LLM on ground robots are provided in [38]. For these experiments, we use the GPT-4o and Gemini-1.5-flash models. All experiments were conducted on a computer with an Intel Core i5 2.4GHz processor and 8 GB of RAM. Software implementation of NAMO-LLM can be found in [37].
IV-A Setting Up Comparative Experiments
NAMO Baselines: We evaluate NAMO-LLM against two baselines: (i) RandomTree, a sampling-based approach from [22] that uses uniform sampling strategies, and (ii) NAMO-SA, a search-based method from [18]. Notably, RandomTree can be recovered from our method by setting $p_{\text{rand}}=p_{\text{obs}}=0$. The approach in [18] selects obstacles for removal by computing the total distance from the robot to the obstacle and from the obstacle to the goal region using the $\text{A}^{*}$ algorithm. This is done for all manipulable objects, and the obstacle with the lowest total distance is selected. The chosen obstacle is then moved to a pre-defined storage room if a collision-free path towards that storage room exists. If no path to the storage room exists, the manipulable obstacle with the next lowest cost is selected; more details can be found in [18]. In our implementation of [18], we model the storage room as an external region adjacent to a specified portion of the workspace boundary. Once an obstacle reaches this boundary segment, it is assumed to be instantaneously transferred to the storage room. On the other hand, our method designs custom drop-off locations for the obstacles during the sampling process. Note that this favors the baseline as our method may sample invalid drop-off locations (e.g., configurations lying in the obstacle space), resulting in sample rejection; see lines 2-2 in Alg. 2. Also in [18], once an obstacle is moved to the storage room, it is never considered for relocation again. This design choice favors that baseline, as the number of obstacles it needs to handle decreases over time—unlike our method. Thus, the maximum horizon $H$ of a plan designed by [18] is equal to the number $n$ of movable obstacles, while in ours it is possible that $H>n$ since our planner is allowed to reconfigure obstacles multiple times. We also note here that [18] includes re-planning strategies to handle unknown obstacles (i.e., obstacles with unknown locations and movability). Since our focus is on known environments, we exclude these capabilities. Finally, to ensure fair comparisons, our implementation of NAMO-LLM ([line 2, Alg. 2]), [18], and [22] share the same RRT planner to check if there exists a collision-free path towards the drop-off location of an obstacle and the storage area.
Robot and Obstacle Configuration Space: We consider a disk-shaped robot with radius $r=0.2$ (Case Studies I-IV) and $r=0.5$ (Case Studies V-VI) and configuration space ${\mathcal{P}}$ determined by the position of the center of the robot. We consider polygonal obstacles with configuration space ${\mathcal{C}}_{i}$ defined by their position and orientation.
Evaluation Metrics: As evaluation metrics, we use (i) the runtime to compute the first feasible plan and (ii) the horizon $H$ of the designed plan. Runtimes and plan horizons $H$ for our method and the sampling-based baseline [22] are averaged over $100$ trials.
[…]
IV-B Comparative Experiments against NAMO Baselines
In what follows, we present six case studies that differ in their complexity in terms of the total number $n$ of movable obstacles, the initial configuration of the obstacles, and the minimum horizon, denoted by $H_{\text{min}}$, of a plan $\tau$ required to reach ${\mathcal{P}}_{g}$. In our method, NAMO-LLM$(p_{\text{rand}},p_{\text{obs}})$, we employ distinct parameter settings for $p_{\text{rand}}$ and $p_{\text{obs}}$. The following results are obtained using the pre-trained GPT 4o model.
Method $|{\mathcal{V}}|$ Time (secs) $H$
NAMO-SA [18] N/A 99.05 27
RandomTree [22] 15.3 14.94 4.3
K=1 NAMO-LLM(0.2, 0.2) 6.6 6.42 3.1
NAMO-LLM(0.2, 0.8) 3.15 3.21 2.5
NAMO-LLM(0.8, 0.2) 6.26 5.74 3.74
NAMO-LLM(0.8, 0.8) 2.71 2.88 2.52
K=2 NAMO-LLM(0.2, 0.2) 7.85 8.77 3.6
NAMO-LLM(0.2, 0.8) 3.30 4.91 2.65
NAMO-LLM(0.8, 0.2) 5.7 4.94 3.5
NAMO-LLM(0.8, 0.8) 3.11 4.46 2.89
K=3 NAMO-LLM(0.2, 0.2) 7.05 7.68 3.10
NAMO-LLM(0.2, 0.8) 3.30 4.47 2.43
NAMO-LLM(0.8, 0.2) 5.45 5.14 3.35
NAMO-LLM(0.8, 0.8) 3.21 4.25 2.67
[…]
Method $|{\mathcal{V}}|$ Time (secs) $H$
[…]
RandomTree [22] 34.90 42.78 4.57
5
NAMO-LLM(0.2, 0.8) 3.86 9.99 2.75
NAMO-LLM(0.8, 0.8) 4.47 7.80 3.04
[…]
RandomTree [22] 92.72 71.32 5.72
5
NAMO-LLM(0.2, 0.8) 6.33 8.52 3.70
NAMO-LLM(0.8, 0.8) 4.37 5.91 3.74
[…]
RandomTree [22] 290.83 408.68 7.41
5
NAMO-LLM(0.2, 0.8) 8.23 19.31 4.16
NAMO-LLM(0.8, 0.8) 5.70 17.12 4.5
[…]
Case Study I ($n=50,H_{\text{min}}=1$): In this case study, we consider a $50\times 50$ environment with $n=50$ movable obstacles. The initial configurations of the obstacles and the robot, as well as the desired robot configuration, are depicted in Fig. 5. The environment is designed such that a plan with horizon $H_{\text{min}}=1$ exists to reach ${\mathcal{P}}_{g}$, meaning that the minimum number of obstacles that must be reconfigured to reach ${\mathcal{P}}_{g}$ is $1$. We test NAMO-LLM for various values of the parameters $p_{\text{rand}}$, $p_{\text{obs}}$, and $K$. The comparative results are summarized in Table I. Observe that our method significantly outperforms both baselines in terms of both the runtime required to compute the first feasible plan and the horizon $H$ of that plan. For instance, the runtimes of the search- and sampling-based methods are $99.05$ and $14.94$ seconds, respectively, while our method can compute the first feasible plan significantly faster; e.g., the average runtime for $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$ is $2.88$ seconds. The average horizon of the plan computed by our method is consistently less than $4$, which is comparable to that of the sampling-based baseline, while the horizon of the search-based one is $27$. We note that the performance of our method does not change significantly with respect to $K\in\{1,2,3\}$. We attribute this to the fact that there are multiple candidate obstacles near ${\mathcal{P}}_{g}$ that can be removed to get access to it and that the employed LLM is capable of identifying them.
Case Study II ($n=100,H_{\text{min}}=1$): In this case study, we consider a $65\times 65$ environment containing $n=100$ movable obstacles, where ${\mathcal{P}}_{g}$ can be reached by removing at least one obstacle; see Fig 5. The comparative results are shown in Table II. Due to space limitations, we present results for our method only with $K=1$. Observe that the performance gap between our method and the baselines increases as $n$ grows. For instance, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, our method finds the first feasible plan in $7.80$ seconds with an average horizon of $H=3.04$. In contrast, [22] takes $42.78$ seconds to compute a plan with a slightly longer horizon ($H=4.57$), while [18] requires $876.36$ seconds and produces a plan with $H=65$.
Case Study III ($n=50,H_{\text{min}}=2$): Here we consider a $50\times 50$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5. The results in Table II show that the performance gap between our method and the baselines further increases as the minimum required horizon $H_{\text{min}}$ grows.
Case Study IV ($n=100,H_{\text{min}}=2$): This case study examines a more complex environment with dimensions $65\times 65$ and $n=100$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5. The results are summarized in Table II, showing that the performance gap further increases as both $n$ and the minimum horizon $H_{\text{min}}$ increase. For instance, our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, finds the first feasible plan in $17.12$ seconds with an average horizon of $H=4.5$. In contrast, the sampling-based approach takes $408.68$ seconds to compute a plan with a longer horizon of $7.41$, while the search-based method requires $877.97$ seconds and produces a plan with $H=65$.
[…]
Case Study V ($n=50,H_{\text{min}}=6$): In this case study, we consider a $40\times 40$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least six obstacles (see Fig. 6). This scenario is significantly more complex due to the larger number of obstacles that need to be relocated. This complexity is also reflected in the increased runtimes reported in Table III, compared to the previous case studies. Our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, computed the first feasible plan in $48.22$ seconds, with an average horizon of $H=26.48$. In contrast, [22] failed to find a feasible plan within $6$ hours while [18] can identify a feasible plan within $141.91$ seconds with $H=42$.
Case Study VI ($n=10,H_{\text{min}}=3$): Finally, we consider a more complex scenario to illustrate the utility of parameter $K$, where—unlike previous case studies—some manipulable obstacles are blocked by others and cannot be relocated until the blocking ones are moved. Specifically, we consider a $20\times 20$ environment with $n=10$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires relocating the manipulable obstacle $o_{1}$. However, unlike Case Studies I-IV, there do not exist collision-free paths to any possible configuration $c_{1}^{\prime}$ unless both $o_{8}$ and $o_{2}$ are moved first; see Fig 6. Thus, in this case, at least $3$ obstacles must be relocated to reach ${\mathcal{P}}_{g}$. The comparative results are shown in Table IV. Unlike Case Study I, the performance of our method improves significantly for $K\in\{2,3\}$ compared to $K=1$. This occurs because, with $K=1$, the LLM often prioritizes removing $o_{1}$ first (i.e., $o_{\text{LLM}}=o_{1}$ in (2)), even though it cannot be moved due to being blocked by $o_{8}$ and $o_{2}$. Consequently, any configuration $c_{1}^{\prime}$ sampled for $o_{1}$ is rejected by line 2 in Alg. 2, slowing down the overall algorithm. However, when $K>1$, the LLM generates a set of obstacles that need to be removed to access ${\mathcal{P}}_{g}$, which may include $o_{2}$ and/or $o_{8}$. Randomly selecting an obstacle from this set helps the algorithm avoid getting stuck due to persistent incorrect LLM recommendations and, consequently, improves performance. Also, observe that our method outperforms [22], which relies on random exploration, in terms of runtime—even when configured with $K=1$. We attribute this to the fact that the LLM occasionally provides correct guidance (i.e., recommending the removal of $o_{2}$ and $o_{8}$ before $o_{1}$) after some redundant actions have been applied (e.g., relocating $o_{3}$ first), as these actions influence the input prompt to the LLM. Overall, our method outperforms both baselines when $K$ is $2$ or $3$. Observe in Table IV that [18] eventually moved all obstacles to the storage area leading to plans with $H=10$.
Summary of Comparisons: Overall, the performance gap between NAMO-LLM and the baselines increases as $n$ or $H_{\text{min}}$ grows. In general, we observed that our method tends to achieve its lowest runtime for computing the first feasible plan when configured with $p_{\text{rand}}=p_{\text{obs}}=0.8$ and $K=1$; if manipulable obstacles are not blocked by other obstacles; see e.g., Case Studies I–V. We attribute this to the LLM’s ability to correctly reason about which obstacle to remove in these scenarios. However, when $p_{\text{rand}}=0.2$ and $p_{\text{obs}}=0.8$, the generated plans tend to have a shorter horizon $H$. This occurs because a smaller $p_{\text{rand}}$ enables the algorithm to select tree nodes for expansion more uniformly, leading to better-quality first plans, albeit with slightly higher runtimes. We also observed that pre-trained LLMs may occasionally make infeasible recommendations regarding which manipulable obstacle to relocate, especially when these obstacles are blocked by others and cannot be repositioned until the blocking ones are moved first; see e.g., Case Study VI. In such cases, setting $K>1$ is beneficial, as it prevents the algorithm from getting stuck due to poor LLM recommendations. Finally, we note that these results depend on the selected LLM.
Method $|{\mathcal{V}}|$ Time (secs) $H$
NAMO-SA [18] N/A 141.91 42
RandomTree [22] over 30k over 22k N/A
NAMO-LLM(0.2, 0.8) 106.12 143.83 19.65
NAMO-LLM(0.8, 0.8) 34.78 48.22 26.48
[…]
Method $|{\mathcal{V}}|$ Time (secs) $H$
NAMO-SA [18] N/A 87.28 10
RandomTree [22] 326.65 227.46 8.95
K=1 NAMO-LLM(0.2, 0.2) 112.45 188.95 10.1
NAMO-LLM(0.2, 0.8) 32.8 359.60 8.11
NAMO-LLM(0.8, 0.2) 85.6 162.09 10.8
NAMO-LLM(0.8, 0.8) 18.45 187.31 8.54
K=2 NAMO-LLM(0.2, 0.2) 82.9 81.92 10.6
NAMO-LLM(0.2, 0.8) 18.5 52.66 7.95
NAMO-LLM(0.8, 0.2) 32.3 37.69 11.1
NAMO-LLM(0.8, 0.8) 12.7 31.05 9.5
K=3 NAMO-LLM(0.2, 0.2) 70.95 98.02 10.65
NAMO-LLM(0.2, 0.8) 26.25 93.51 8.35
NAMO-LLM(0.8, 0.2) 34.9 44.23 12.05
NAMO-LLM(0.8, 0.8) 12.65 42.47 9.9
[…]
Method $|{\mathcal{V}}|$ Time (secs) $H$
[…]
RandomTree [22] 290.83 408.68 7.41
NAMO-LLM(0.2, 0.8) 84.35 233.94 15.8
NAMO-LLM(0.8, 0.8) 24.75 104.30 17.2
GPT 4o 5.70 17.12 4.16
[…]
RandomTree [22] over 30k over 22k N/A
NAMO-LLM(0.2, 0.8) 27.70 21.03 9.6
NAMO-LLM(0.8, 0.8) 15.15 14.24 12.15
GPT 4o 34.78 48.22 19.65
TABLE V: Effect of LLM Model on average tree size, runtimes, and horizon. The line ‘GPT 4o’ reports the best results of our method when paired with GPT 4o.
IV-C Comparative Results with an Alternative LLM
In this section, we demonstrate the performance of our method when paired with the Gemini-1.5-flash model on Case Studies IV ($n=100$, $H_{\text{min}}=2$) and V ($n=50$, $H_{\text{min}}=6$). The comparative results are shown in Table V. In Case Study IV, our method outperforms both baselines in terms of runtime but produces longer plans than the sampling-based baseline (while still shorter than those of the search-based one). In Case Study V, however, our approach achieves better performance than both baselines in terms of both runtime and plan horizon $H$. Overall, we observe that switching from GPT-4o to Gemini-1.5-flash tends to improve the performance of our algorithm in scenarios where manipulable obstacles are blocked by other obstacles (see Case Study V) and deteriorate its performance in cases with large number of movable obstacles (see Case Study IV). Nevertheless, our method still remains more scalable than the baselines.
V Conclusion
In this paper, we proposed NAMO-LLM, a probabilistically complete sampling-based planner for NAMO problems. Through comparative experiments, we demonstrated its scalability and efficiency in highly cluttered environments. Future work will focus on extending NAMO-LLM to settings with obstacles of unknown location and movability.

### v2

@@
 To address this, we propose NAMO-LLM, a sampling-based planner that searches over robot and obstacle configurations to compute feasible plans specifying which obstacles to move, where, and in what order.
-Its key novelty is a non-uniform sampling strategy guided by Large Language Models (LLMs) biasing the tree construction toward directions more likely to yield a solution.
+Its key novelty is a non-uniform sampling strategy guided by Large Language Models (LLMs) biasing the search process toward directions more likely to yield a solution.
 We show that NAMO-LLM is probabilistically complete and demonstrate through experiments that it efficiently scales to cluttered environments, outperforming related works in both runtime and plan quality.
 I Introduction
-Several planning algorithms have been developed to compute robot paths from an initial configuration to a desired one while avoiding obstacles [1, 2, 3, 4, 5].
-However, in highly cluttered environments, obstacle-free paths to goal regions may not exist.
-In such cases, rather than reporting task failure, a more effective approach is to enable robots to interact with their environment by repositioning movable obstacles.
-This requires planning algorithms capable of determining which obstacles to move, in what order, and where to reposition them to create a feasible path to the goal; see Fig.
+Several planning algorithms have been developed to compute robot paths from an initial configuration to a desired one while avoiding obstacles [1, 2].
+However, in highly cluttered environments, obstacle-free paths to goal regions may not exist; see Fig.
 1.
-This planning problem, known as Navigation Among Movable Objects (NAMO) [6], has been proven to be NP-hard [7].
-Several algorithms have been proposed to address NAMO problems.
-Early approaches relied on heuristic solutions which lacked completeness guarantees [8].
+In such cases, rather than reporting task failure, a more effective approach is to enable robots to interact with their environment by relocating movable obstacles.
+This requires planners capable of determining which obstacles to move, in what order, and where to reposition them to create a feasible path to the goal; see Fig.
+1.
+This problem, known as Navigation Among Movable Objects (NAMO) has been proven to be NP-hard [3].
+Early approaches to address NAMO problems relied on heuristic solutions which lacked completeness guarantees [4].
 To overcome this, search-based planning methods have been introduced.
 These methods either provide completeness guarantees for specific subclasses of NAMO problems or sacrifice computational efficiency to ensure completeness across a broader range of scenarios [5, 6, 7, 8].
-For instance, [6] addressed NAMO problems where disconnected free-space components can be connected by moving a single obstacle at a time.
-This formulation allows the problem to be represented as a graph structure, enabling the use of resolution-complete planners.
-This work was later extended to handle cases where at most $k$ obstacles must be moved, with each obstacle requiring only a single displacement [10].
-A more general approach is proposed in [11] to solve non-monotone NAMO problems, where multiple obstacles may need to be repositioned—potentially more than once—to establish connectivity between disconnected free-space components.
-While this algorithm is complete under mild assumptions, its computational cost scales exponentially with the number of movable objects.
-Extensions of these approaches to account for physical interactions between objects are presented in [12, 13], while extensions to unknown environments are considered in [14, 15, 16, 17, 18, 19].
-In addition to search-based methods, sampling-based planners have been developed to improve efficiency while still maintaining (probabilistic) completeness guarantees for general NAMO problems [20, 21, 22, 23].
-The key idea in these approaches is to incrementally construct trees by sampling and adding nodes, where each node represents the environment’s configuration, and edges model feasible manipulation actions.
-However, their efficiency declines as the number of movable objects or the number of required obstacle relocations grows, due to their uniform sampling strategy.
-Figure 1: A ground robot is tasked with reaching a destination behind the building block; see (1).
-All direct pathways are initially blocked (see Fig.
-2).
-The robot autonomously removes obstacles to clear a path; see (2)–(4).
-This paper addresses this limitation by designing a new sampling-based NAMO algorithm that efficiently scales to environments characterized by large numbers of movable obstacles and/or the need for multiple obstacle relocations to reach desired regions.
-Inspired by [22], our algorithm incrementally builds trees that explore the space of obstacle and robot configurations.
-Each tree node represents an environment configuration, while edges among nodes correspond to manipulation actions that relocate movable obstacles.
-At each iteration, a tree node and a manipulation action are sampled.
-Application of that action to the sampled tree node results in new tree node that is added to the tree if feasible.
-The key novelty of our work lies in a non-uniform sampling strategy that leverages Large Language Models (LLMs) to accelerate plan generation.
-Specifically, our sampling strategy directs the tree to grow towards LLM-recommended directions at certain times while allowing random exploration otherwise.
-We refer to the resulting planner as NAMO-LLM, for Navigation Among Movable Obstacles with Large Language Model guidance.
-We show that NAMO-LLM is probabilistically complete meaning that the probability of computing a feasible plan converges to one as the number of iterations goes to infinity.
-We conduct extensive comparative experiments against both search-based planners [18] and sampling-based planners relying on uniform sampling strategies [22], demonstrating that our method significantly reduces both the time required to compute the first feasible plan and the corresponding number of obstacles that need to be relocated.
-We also demonstrate NAMO-LLM on hardware experiments; see Fig.
-1.
-Related Works: In addition to the NAMO algorithms discussed earlier, our work is also related to research on rearrangement planning (RP) problems where robots manipulate movable obstacles [24, 25, 26, 27].
+Extensions of these approaches to unknown environments are considered in [9, 10, 11, 12, 13].
+Sampling-based planners have been developed to improve efficiency while still maintaining (probabilistic) completeness guarantees for general NAMO problems [14, 15, 16, 17].
+Their key idea is to incrementally construct trees by sampling nodes, where each node represents the environment’s configuration, and edges model feasible manipulation actions.
+However, their efficiency declines as the number of movable obstacles or the number of required obstacle relocations grows, due to their uniform sampling strategy.
+Figure 1: A robot must reach a destination behind the building block. Since all direct paths are blocked, it relocates obstacles to access the goal region.
+This paper addresses this limitation by designing a new sampling-based NAMO algorithm that efficiently scales to environments with many movable obstacles and/or requiring multiple obstacle relocations to reach target regions.
+Inspired by [16], our algorithm incrementally builds trees exploring the configuration space of movable objects and the robot’s free-space components.
+The key novelty lies in a non-uniform sampling strategy that leverages Large Language Models (LLMs) to accelerate plan generation.
+Specifically, the strategy directs the tree toward LLM-recommended directions at certain times while allowing random exploration otherwise.
+We refer to the resulting planner as NAMO-LLM and show it is probabilistically complete, meaning the probability of computing a feasible plan (if one exists) converges to one as the number of iterations grows.
+Extensive numerical and hardware experiments demonstrate that NAMO-LLM outperforms search-based planners [12], uniform-sampling planners [16], and planners relying solely on LLMs in terms of runtime for the first feasible plan, the number of obstacles relocated, and task success rates.
+This performance gap increases with environment complexity.
+Related Works: In addition to the NAMO algorithms discussed earlier, our work is also related to rearrangement planning (RP) problems where robots manipulate movable obstacles [18, 19, 20, 21].
 RP differs from NAMO because it considers predefined goal positions for the movable obstacles and typically does not specify a final goal for the robot.
-While our work primarily focuses on NAMO problems, we believe that our non-uniform sampling strategy can also be integrated with various sampling-based approaches for planning among movable obstacles, including RP and manipulation planning [28, 29].
-However, a formal integration with these approaches is beyond the scope of this paper.
-Furthermore, several works design non-uniform sampling strategies to accelerate sampling-based planners [30, 31, 32, 33, 34, 35, 36].
-For example, [34] learns a non-uniform sampling distribution from demonstrations.
+We believe that our non-uniform sampling strategy can also be integrated with various sampling-based approaches for planning among movable obstacles, including RP and manipulation planning [22, 23]; however, this integration is beyond the scope of this paper.
+Furthermore, several works design non-uniform sampling strategies to accelerate sampling-based planners [24, 25, 26, 27, 28].
 However, we emphasize that these works target different planning problems from ours, focusing on designing obstacle-free paths between an initial and a desired robot configuration (while also assuming that such paths exist).
 As such, they cannot be directly applied to our setting.
-It is worth noting that, unlike [32, 33, 34, 35, 36], our work does not require training neural networks; instead, ours leverages pre-trained LLMs.
+It is worth noting that, unlike [25, 26, 27, 28], our work does not require training neural networks; instead, ours leverages pre-trained LLMs.
 Summary of Contributions: First, we introduce NAMO-LLM, a novel sampling-based planner designed to efficiently solve NAMO problems in highly cluttered environments.
-Second, to the best of our knowledge, this is the first integration of LLMs within a sampling-based NAMO planner, significantly accelerating plan generation.
-Third, we conduct extensive experiments demonstrating that our approach outperforms existing planning algorithms in terms of efficiency and solution quality.
-Fourth, we release an open-source implementation of our proposed algorithm [37].
+Second, we propose the first integration of LLMs within a sampling-based NAMO planner, significantly accelerating plan generation.
+Third, we conduct extensive experiments demonstrating that our approach scales well to complex environments and outperforms related planners in efficiency, solution quality, and task success rates.
 II Problem Formulation
-Robot Modeling: We consider a robot with the following kinematics: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ where $\mathbf{p}(t)\in{\mathcal{P}}$ and $\mathbf{u}(t)\in{\mathcal{U}}$ denote the robot configuration (e.g., position and orientation) and the control input at time step $t$, respectively.
-We assume that ${\mathbf{p}}(t)$ is known at all times and that the robot is holonomic and, therefore, it can move along any desired path.
+Robot Modeling: We consider a robot modeled as: ${\mathbf{p}}(t+1)={\mathbf{f}}({\mathbf{p}}(t),{\mathbf{u}}(t)),$ where $\mathbf{p}(t)\in{\mathcal{P}}$ and $\mathbf{u}(t)\in{\mathcal{U}}$ denote the robot configuration (e.g., position and orientation) and the control input at time step $t$, respectively.
+We assume that ${\mathbf{p}}(t)$ is always known and that the robot is holonomic and, therefore, it can move along any desired path.
 The robot is also equipped with a gripper, enabling it to manipulate one movable object at a time.
-Figure 2: Top-view representation of the urban environment shown in Fig.
-1.
-The set ${\mathcal{P}}$ is defined by the position of the center of the disk-shaped robot.
-Given ${\mathcal{N}}(0)$, the set ${\mathcal{O}}_{M}(0)$ includes only obstacles 1-6.
-Environment Modeling: The robot operates in an environment $\Omega$ containing a set of known obstacles ${\mathcal{O}}=\{o_{1},\dots,o_{O}\}$.
+Environment Modeling: The robot resides in an environment $\Omega$ containing a set of known obstacles ${\mathcal{O}}=\{o_{1},\dots,o_{O}\}$.
 The obstacles are classified as movable obstacles, collected in a set ${\mathcal{O}}_{\text{movable}}\subseteq{\mathcal{O}}$, and non-movable ones.
@@
 The configuration (e.g., position and orientation) of an obstacle $o_{i}\in{\mathcal{O}}$ at time $t$ is denoted by $c_{i}(t)\in{\mathcal{C}}_{i}$ where ${\mathcal{C}}_{i}$ denotes its configuration space.
-We also denote by $co_{i}(t)$ the region in the configuration space $\mathcal{P}$ of the robot that corresponds to configurations where the robot is in collision with obstacle $o_{i}$; see Fig.
+We also denote by ${\bf{co}}_{i}(t)$ the region in $\mathcal{P}$ that corresponds to configurations where the robot is in collision with obstacle $o_{i}$.
+Thus, the robot’s free space at time $t$, denoted by ${\mathcal{P}}_{\text{free}}(t)$, is defined as ${\mathcal{P}}_{\text{free}}(t)={\mathcal{P}}\setminus\mathcal{CO}(t)$ where $\mathcal{CO}(t)=\cup_{i=1}^{n}\{{\bf{co}}_{i}(t)\}$.
+Notice that ${\mathcal{P}}_{\text{free}}(t)$ may consist of multiple connected components whose boundaries are determined by the configuration of the obstacles.
+We denote by ${\mathcal{N}}(t)\subseteq{\mathcal{P}}_{\text{free}}(t)$, the free-space component that ${\mathbf{p}}(t)$ lies in; see also Fig.
 2.
-Thus, the robot’s free space at time $t$, denoted by ${\mathcal{P}}_{\text{free}}(t)$, is defined as ${\mathcal{P}}_{\text{free}}(t)={\mathcal{P}}\setminus\mathcal{CO}(t)$ where $\mathcal{CO}(t)=\cup_{i=1}^{n_{O}}\{co_{i}(t)\}$.
-Notice that ${\mathcal{P}}_{\text{free}}(t)$ may consist of multiple connected components whose boundaries are determined by the configuration of the obstacles.
-We denote by ${\mathcal{N}}(t)\subseteq{\mathcal{P}}_{\text{free}}(t)$, the free-space component that ${\mathbf{p}}(t)$ lies in; see Fig.
-2.
-Robot Task: The robot is tasked with reaching a desired configuration starting from an initial one ${\mathbf{p}}(0)$ while avoiding collisions with obstacles.
-Hereafter, we denote by ${\mathcal{P}}_{g}\subseteq{\mathcal{P}}$ a set collecting all desired robot configurations.
-We consider cases where all configurations in ${\mathcal{P}}_{g}$ are initially unreachable due to the initial configuration of obstacles, i.e., ${\mathcal{P}}_{g}\cap{\mathcal{N}}(0)=\emptyset$.
-In such cases, the robot must determine which movable obstacles need to be relocated and where they should be placed.
-To formalize this, we define the set of manipulable obstacles at time $t$ as $\mathcal{O}_{M}(t)\subseteq{\mathcal{O}}_{\text{movable}}$ which includes all movable obstacles $o_{i}$ that the robot can directly reach, i.e., $\mathcal{O}_{M}(t)=\{o_{i}\in{\mathcal{O}}_{\text{movable}}\mid co_{i}(t)\cap{\mathcal{N}}(t)\neq\emptyset\}$; see also Fig.
-2.
-Next, we define manipulation actions of the form $a(c_{i}(t),c_{i}^{\prime})$ which involve moving a manipulable object $o_{i}\in\mathcal{O}_{M}(t)$, from its current configuration $c_{i}(t)$ to a new configuration $c_{i}(t^{\prime})=c_{i}^{\prime}$, for some $t^{\prime}>t$.
-In other words, the application of this action requires the robot to approach the obstacle’s current location (determined by $c_{i}(t)$), grasp it, and relocate it to the new configuration $c_{i}^{\prime}$.
-Thus, our goal is to compute a plan $\tau$, i.e., a sequence of manipulation actions so that eventually ${\mathcal{P}}_{g}\cap{\mathcal{N}}(t)\neq\emptyset$.
-We defined this plan as $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, where $\tau(k)$ refers to the $k$-th manipulation action, $k\in\{0,1,\dots,H\}$, for some horizon $H\geq 0$.
-Note that the index $k$ is different from the time step $t$, as the application of an action $\tau(k)$ may require multiple time steps $t$.
-The problem that this paper addresses can be summarized as follows:
+Robot Task: The robot must reach a desired configuration from an initial one ${\mathbf{p}}(0)$ while avoiding obstacles.
+Let ${\mathcal{P}}_{g}\subseteq{\mathcal{P}}$ denote the desired configurations.
+We focus on cases where all configurations in ${\mathcal{P}}_{g}$ are initially unreachable due to obstacle placements, i.e., ${\mathcal{P}}_{g}\cap{\mathcal{N}}(0)=\emptyset$.
+The robot must then determine which movable obstacles to relocate and their target placements.
+To formalize this, we define the set of manipulable obstacles at time $t$ as $\mathcal{O}_{M}(t)\subseteq{\mathcal{O}}_{\text{movable}}$ which includes all movable obstacles $o_{i}$ that the robot can directly reach, i.e., $\mathcal{O}_{M}(t)=\{o_{i}\in{\mathcal{O}}_{\text{movable}}\mid{\bf{co}}_{i}(t)\cap{\mathcal{N}}(t)\neq\emptyset\}$.
+Next, we define manipulation actions of the form $a(c_{i}(t),c_{i}^{\prime})$ requiring the robot to approach, grasp, and relocate an an object $o_{i}\in\mathcal{O}_{M}(t)$, from its current configuration $c_{i}(t)$ to a new configuration $c_{i}(t^{\prime})=c_{i}^{\prime}$, for some $t^{\prime}>t$.
+Our goal is to compute a plan $\tau=\tau(0),\dots,\tau(H)$ such that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$ for some horizon $H\geq 0$.
+Here, $\tau(k)$ denotes the $k$-th manipulation action, with $k\in\{0,1,\dots,H\}$.
+Note that $k$ indexes actions, not time steps, as executing $\tau(k)$ may span multiple time steps $t$.
 Problem 1
-Given an initial robot configuration ${\mathbf{p}}(0)$, a set ${\mathcal{P}}_{g}$ of desired configurations, and initial obstacle configurations $c_{i}(0)$, compute a manipulation plan $\tau=\tau(0),\tau(1),\dots,\tau(k),\dots,\tau(H)$, so that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$.
+Given ${\mathbf{p}}(0)$, ${\mathcal{P}}_{g}$, $\mathcal{O}$}, ${\mathcal{O}}_{\text{movable}}$, and $c_{i}(0)$ for all $o_{i}$, compute $\tau=\tau(0),\dots,\tau(H)$ so that ${\mathcal{P}}_{g}\cap{\mathcal{N}}(H)\neq\emptyset$.
 III Proposed Planning Algorithm
-In this section, we present NAMO-LLM, our proposed planner to address Problem 1.
-In Section III-A, we present a sampling-based planner to design plans $\tau$.
-In Section III-B, we present a non-uniform sampling strategy that accelerates the plan synthesis phase by leveraging Large Language Models (LLM).
+In this section, we present NAMO-LLM, our planner for Problem 1.
+Section III-A describes a sampling-based planner for generating plans $\tau$.
+Section III-B introduces a non-uniform sampling strategy leveraging LLMs to accelerate plan synthesis.
 In Section III-C, we discuss the input prompt to the LLM guiding the sampling strategy.
 In Section III-D, we show that NAMO-LLM is probabilistically complete.
+Figure 2: Tree structure built by the proposed planner.
+Each node (blue) represents a state ${\mathbf{q}}(t)$, and each edge (green arrow) represents a manipulation action.
+For instance, ${\mathbf{q}}_{\text{new}}$ is reached from ${\mathbf{q}}_{\text{rand}}$ via action $a({\mathbf{c}}_{5},{\mathbf{c}}_{5}^{\prime})$ moving obstacle $o_{5}$.
+If moving $o_{5}$ is recommended by the LLM (rather than random sampling), the score $v({\mathbf{q}}_{\text{new}})$ is set to 11; otherwise, it remains 0.
+The label next to each node indicates whether it was generated by moving an LLM-suggested obstacle (L) or a randomly selected obstacle (R) (see Section III-B).
+The number next to the label is the score $u({\mathbf{q}})$ yielding a set ${\mathcal{V}}_{v}$ that collects the top two gray-shaded nodes (see Section III-B).
+The two snapshots illustrate the environment associated with ${\mathbf{q}}_{\text{rand}}$ and ${\mathbf{q}}_{\text{new}}$.
+III-A Sampling-based Planning
+We propose a sampling-based planner to address Problem 1, summarized in Alg.
+1.
+The algorithm incrementally builds a tree exploring the configuration space of the movable objects and the robot’s free-space components to compute $\tau$; see Fig.
+2.
+We denote the constructed tree by ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where $\mathcal{V}$} and $\mathcal{E}$} is a set of nodes and edges, respectively.
+The set $\mathcal{V}$} contains states of the form ${\mathbf{q}}(t)=[\mathcal{N}(t),{\mathbf{c}}(t)]$, where ${\mathbf{c}}(t)$ is a vector stacking the configuration of all obstacles, i.e., ${\mathbf{c}}(t)=[{\mathbf{c}}_{1}(t),{\mathbf{c}}_{2}(t),\dots,{\mathbf{c}}_{n}(t)]$ and ${\mathcal{N}}(t)$ denotes the free-space component that ${\mathbf{p}}(t)$ lies in.
+The root of the tree models the initial configuration of the obstacles and the robot, i.e., ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ [lines 1-1, Alg.
+1].
+The tree is initialized so that ${\mathcal{V}}=\{{\mathbf{q}}(0)\}$ and $\mathcal{E}$$\emptyset$ [line 1, Alg.
+1].
+Also, the tree is built incrementally by sampling new states ${\mathbf{q}}_{\text{new}}$ that are added to the tree structure [line 1, Alg.
+1].
+The algorithm terminates after a user-specified number of iterations $m_{\text{max}}$ and returns a feasible plan $\tau$ (if it has been found).
+In what follows, we present our method in detail.
 Input: (i) initial robot configuration ${\mathbf{p}}(0)$; (ii) initial obstacle configuration ${\mathbf{c}}(0)$; (iii) Set ${\mathcal{O}}_{\text{movable}}$ ; (iv) goal set ${\mathcal{X}}_{g}$
@@
 5  Sample a node ${\mathbf{q}}_{\text{rand}}\in{\mathcal{V}}$ using $f_{\mathcal{V}}$;
-6       $[{\mathbf{q}}_{\text{new}},a(c_{i}(t),c_{i}^{\prime})]=\texttt{TreeExpansion}({\mathbf{q}}_{\text{rand}})$;
+6  ${\mathbf{q}}_{\text{new}},a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})]=\texttt{TreeExpansion}({\mathbf{q}}_{\text{rand}})$;
 7  if ${\mathbf{q}}_{\text{new}}\neq\varnothing$ then
 8     ${\mathcal{V}}\leftarrow{\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$, ${\mathcal{E}}\leftarrow{\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$;
-9             $A(({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}}))=a(c_{i}(t),c_{i}^{\prime})$;
+9     $A(({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}}))=a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})$;
 10     Extract ${\mathcal{N}}(t^{\prime})$ from ${\mathbf{q}}_{\text{new}}=[\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$;
@@
 19  Using $A$ and $\tau_{{\mathbf{q}}}$, compute the manipulation plan $\tau$ ;
-20
-21else
+20  Compute robot motion path $\xi$ executing $\tau$;
+21
+22else
 23  No plan $\tau$ has been found;
-23
+24
 Algorithm 1 NAMO-LLM
-1 Compute the set of manipulable obstacles ${\mathcal{O}}_{M,\text{rand}}(t)$ ;
+1Compute the set of manipulable obstacles ${\mathcal{O}}_{M,\text{rand}}(t)$ ;
 2 Sample a manipulable obstacle $o_{i}$ from a given distribution $f_{{\mathcal{O}}_{M}}$ ;
-3 Sample a random configuration $c_{i}^{\prime}$ from $f_{{\mathcal{C}}_{i}}$;
-4 if $co_{i}^{\prime}\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ and $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$  then
-5       if $c_{i}(t)\to c^{\prime}_{i}$ is collision-free  then
-6             Initialize ${\mathbf{c}}(t^{\prime})={\mathbf{c}}(t)$ and update $c_{i}(t^{\prime})=c_{i}^{\prime}$ in ${\mathbf{c}}(t^{\prime})$;
-7             $\mathcal{P}_{\text{free}}(t^{\prime})=\mathcal{P}_{\text{free}}(t)\cup co_{i}(t)\setminus co_{i}^{\prime}$ ;
-8             Given ${\mathbf{c}}(t^{\prime})$, select $\mathcal{N}(t^{\prime})$ ;
-9             ${\mathbf{q}}_{\text{new}}=(\mathcal{N}(t^{\prime}),{\mathbf{c}}(t^{\prime}))$ ;
-10             return ${\mathbf{q}}_{\text{new}}$, $a(c_{i}(t),c_{i}^{\prime})$
+3 Sample a random configuration ${\mathbf{c}}_{i}^{\prime}$ from $f_{{\mathcal{C}}_{i}}$;
+4 if ${\bf{co}}_{i}^{\prime}\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ and ${\bf{co}}_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}(t)\setminus{\bf{co}}_{i}(t)$ then
+5  Select ${\mathbf{p}}(t)\in{\mathcal{N}}_{\text{rand}}(t)$ and $\mathcal{N}$$^{\prime}$ and compute $\zeta$;
+6  if $\exists~\zeta$ then
+7     Initialize ${\mathbf{c}}(t^{\prime})={\mathbf{c}}(t)$ and update ${\mathbf{c}}_{i}(t^{\prime})={\mathbf{c}}_{i}^{\prime}$ in ${\mathbf{c}}(t^{\prime})$;
+8     $\mathcal{P}_{\text{free}}(t^{\prime})=\mathcal{P}_{\text{free}}(t)\cup{\bf{co}}_{i}(t)\setminus{\bf{co}}_{i}^{\prime}$ ;
+9     ${\mathbf{q}}_{\text{new}}=(\mathcal{N}^{\prime},{\mathbf{c}}(t^{\prime}))$ ;
+10     return ${\mathbf{q}}_{\text{new}}$, $a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})$
 11 else
-12             return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
+12     return ${\mathbf{q}}_{\text{new}}=\varnothing,~a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})=\varnothing$
 13else
-14       return ${\mathbf{q}}_{\text{new}}=\varnothing,~{}a(c_{i}(t),c_{i}^{\prime})=\varnothing$
+14  return ${\mathbf{q}}_{\text{new}}=\varnothing,~a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})=\varnothing$
 Algorithm 2 TreeExpansion(${\mathbf{q}}_{\text{rand}}$)
-III-A Sampling-based Planning
-We propose a sampling-based planner to address Problem 1 summarized in Alg.
-1 The proposed algorithm incrementally builds a tree that explores the configuration space of the movable objects and the robot’s free space components to compute a manipulation plan $\tau$ as defined in Section II.
-We denote the constructed tree by ${\mathcal{G}}=\{{\mathcal{V}},{\mathcal{E}}\}$, where ${\mathcal{V}}$ and ${\mathcal{E}}$ is a set of nodes and edges, respectively.
-The set ${\mathcal{V}}$ contains states of the form ${\mathbf{q}}(t)=[\mathcal{N}(t),{\mathbf{c}}(t)]$, where ${\mathbf{c}}(t)$ is a vector stacking the configuration of all obstacles, i.e., ${\mathbf{c}}(t)=[c_{1}(t),c_{2}(t),\dots,c_{n}(t)]$ and ${\mathcal{N}}(t)$ denotes the free-space component that ${\mathbf{p}}(t)$ lies in.
-The root of the tree models the initial configuration of the obstacles and the robot, i.e., ${\mathbf{q}}(0)=[{\mathcal{N}}(0),{\mathbf{c}}(0)]$ [lines 1-1, Alg.
-1].
-The tree is initialized so that ${\mathcal{V}}=\{{\mathbf{q}}(0)\}$ and ${\mathcal{E}}=\emptyset$ [line 1, Alg.
-1].
-Also, the tree is built incrementally by sampling new states ${\mathbf{q}}_{\text{new}}$ that are added to the tree structure [line 1, Alg.
-1].
-The algorithm terminates after a user-specified number of iterations $m_{\text{max}}$ and returns a feasible plan $\tau$ (if it has been found).
-In what follows, we present a detailed description of our method.
 At every iteration $m$ of Alg.
@@
 1].
-The mass function $f_{{\mathcal{V}}}$ defines the probability of selecting a node ${\mathbf{q}}\in{\mathcal{V}}$.
 Second, we sample a new node ${\mathbf{q}}_{\text{new}}$ that will be added to the tree; see Alg.
@@
 2].
-The mass function $f_{{\mathcal{O}}_{M}}$ defines the probability of selecting a manipulable obstacle $o_{i}\in{\mathcal{O}}_{M,\text{rand}}$.
-Given the selected manipulable obstacle $o_{i}$, we sample a new configuration $c_{i}^{\prime}\in{\mathcal{C}}_{i}$ using a density function $f_{\mathcal{C}_{i}}:\mathcal{C}_{i}\rightarrow[0,\infty)$ [line 2, Alg.
+Given the selected manipulable obstacle $o_{i}$, we sample a new configuration ${\mathbf{c}}_{i}^{\prime}\in{\mathcal{C}}_{i}$ using a density function $f_{\mathcal{C}_{i}}:\mathcal{C}_{i}\rightarrow[0,\infty)$ [line 2, Alg.
 2].
 Designing the functions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ is deferred to Section III-B.
-Figure 3: Illustration of the free-space components ${\mathcal{N}}(t)$ (left column) and ${\mathcal{N}}^{1}(t)$ (right column) in two environments.
-Next, we check if a collision-free path exists for the robot to move $o_{i}$ from $c_{i}(t)$ to $c_{i}^{\prime}$ [line 2, Alg.
+Next, we check if a collision-free path exists for the robot to execute $a({\mathbf{c}}_{i},{\mathbf{c}}_{i}^{\prime})$, i.e., to move $o_{i}$ from ${\mathbf{c}}_{i}(t)$ to ${\mathbf{c}}_{i}^{\prime}$ [line 2, Alg.
 2].
-Two necessary conditions must be met for $c_{i}^{\prime}$ to admit such a collision-free path; if they are not satisfied then such a path does not exist [line 2, Alg.
+Transporting $o_{i}$ first requires the robot to move from its current configuration in ${\mathcal{N}}_{\text{rand}}$ to ${\bf{co}}_{i}(t)$ in order to grasp $o_{i}$.
+Such a path always exists regardless of the exact robot configuration in ${\mathcal{N}}_{\text{rand}}$ because $o_{i}$ is manipulable (i.e., ${\bf{co}}_{i}\cap{\mathcal{N}}_{\text{rand}}\neq\emptyset$) and the robot is holonomic (see Section II).
+Therefore, in what follows, we focus only on the existence of a collision-free path that enables the robot to move $o_{i}$ from ${\mathbf{c}}_{i}(t)$ to ${\mathbf{c}}_{i}^{\prime}$, assuming the robot is already grasping $o_{i}$.
+Two necessary conditions must be met for ${\mathbf{c}}_{i}^{\prime}$ to admit such a path [line 2, Alg.
 2].
-First, $c_{i}^{\prime}$ must not intersect other obstacles, i.e., $co_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}^{\text{rand}}(t)\setminus co_{i}(t)$, where ${\mathcal{P}}_{\text{free}}^{\text{rand}}(t)$ is computed using ${\mathbf{q}}_{\text{rand}}$.
-Second, $co_{i}^{\prime}$ must not lie in a component of ${\mathcal{P}}_{\text{free}}$ that becomes accessible only if an obstacle $o_{j}\neq o_{i}$ is relocated first.
-To formalize this, we define ${\mathcal{P}}^{i}_{\text{free}}(t)$ as the robot’s free space if obstacle $o_{i}$ is removed from the environment, i.e., ${\mathcal{P}}^{i}_{\text{free}}(t)={\mathcal{P}}_{\text{free}}(t)\setminus co_{i}(t)$.
+First, ${\mathbf{c}}_{i}^{\prime}$ must not intersect other obstacles, i.e., ${\bf{co}}_{i}^{\prime}\in{\mathcal{P}}_{\text{free}}^{\text{rand}}(t)\setminus{\bf{co}}_{i}(t)$, where ${\mathcal{P}}_{\text{free}}^{\text{rand}}(t)$ is computed using ${\mathbf{q}}_{\text{rand}}$.
+Second, ${\bf{co}}_{i}^{\prime}$ must not lie in a component of ${\mathcal{P}}_{\text{free}}^{\text{rand}}$ that becomes accessible only if an obstacle $o_{j}\neq o_{i}$ is relocated first.
+To formalize this, we define ${\mathcal{P}}^{i}_{\text{free}}(t)$ as the robot’s free space if obstacle $o_{i}$ is removed from the environment, i.e., ${\mathcal{P}}^{i}_{\text{free}}(t)={\mathcal{P}}_{\text{free}}(t)\setminus{\bf{co}}_{i}(t)$.
 Correspondingly, ${\mathcal{N}}^{i}(t)$ denotes the component of ${\mathcal{P}}^{i}_{\text{free}}(t)$ containing the robot’s position ${\mathbf{p}}(t)$; see Fig.
 3.
-Thus, the second condition requires $co_{i}^{\prime}$ to belong to ${\mathcal{N}}^{i}_{\text{rand}}(t)$ (i.e., to ${\mathcal{N}}^{i}(t)$ computed using ${\mathbf{q}}_{\text{rand}}$).
-If the above two conditions hold, then there may exist such a collision-free path for $o_{i}$.
-This path is defined as a sequence of obstacle configurations, i.e., $\boldsymbol{\zeta}=c_{i}(t),c_{i}(t+1),\dots,c_{i}(t+s),\dots,c_{i}(t^{\prime})$ such that (i) $c_{i}(t^{\prime})=c_{i}^{\prime}$ and (ii) $co_{i}(t+s)\in{\mathcal{N}}^{i}(t)$, for all $s\in\{1,2,\dots,t^{\prime}-t\}$.
-This can be checked using any existing motion planner such as RRT or A* [3].
-If such a path $\boldsymbol{\zeta}$ exists, this means the obstacle $o_{i}$ can be moved to $c_{i}^{\prime}$ giving rise to ${\mathbf{c}}(t^{\prime})$ that differs from ${\mathbf{c}}(t)$ only in the configuration of the obstacle $o_{i}$ [line 2, Alg.
-2].
-In general, ${\mathcal{N}}^{i}(t)$ may split into multiple components once $o_{i}$ is moved to $c_{i}^{\prime}$, each adjacent to $co_{i}^{\prime}$.
-The robot can be in any of these components, and we randomly select one to be ${\mathcal{N}}(t^{\prime})$ [line 2, Alg.
-2].
-Given the selected free-space component ${\mathcal{N}}(t^{\prime})$ and ${\mathbf{c}}(t^{\prime})$, we construct the node ${\mathbf{q}}_{\text{new}}=[{\mathcal{N}}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$ [line 2, Alg.
-2] and the sets of nodes and edges are updated as ${\mathcal{V}}={\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$ and ${\mathcal{E}}={\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$, respectively [lines 1, Alg.
+Thus, the second condition requires ${\bf{co}}_{i}^{\prime}$ to belong to ${\mathcal{N}}^{i}_{\text{rand}}(t)$ (i.e., to ${\mathcal{N}}^{i}(t)$ computed using ${\mathbf{q}}_{\text{rand}}$).
+Figure 3: Illustration of the components ${\mathcal{N}}(t)$ (left) and ${\mathcal{N}}^{5}(t)$ (right).
+Uniform vs non-uniform sampling for obstacle selection: In the left figure, only obstacles 1, 3, 4, and 5 are manipulable.
+Under uniform sampling, each obstacle $o_{i}$ will be selected with probability $p_{i}$ defined as $p_{1}=p_{3}=p_{4}=p_{5}=0.25$ and $p_{2}=0$.
+Under non-uniform sampling, if the LLM-recommended obstacle $o_{\text{LLM}}$ is $o_{3}$ and $p_{\text{obs}}=0.8$, then the corresponding probabilities become $p_{3}=0.85$, $p_{1}=p_{4}=p_{5}=0.05$, and $p_{2}=0$.
+If these two conditions hold, then a collision-free path to execute $a({\mathbf{c}}_{i},{\mathbf{c}}_{i}^{\prime})$ (as soon as it grasps $o_{i}$) may exist.
+This path is defined for the robot while carrying $o_{i}$.
+Thus, with slight abuse of notation, we define this path as $\zeta=[{\mathbf{p}}(t),{\mathbf{c}}_{i}(t)],[{\mathbf{p}}(t+1),{\mathbf{c}}_{i}(t+1)],\dots,[{\mathbf{p}}(t+s),{\mathbf{c}}_{i}(t+s)],\dots,[{\mathbf{p}}(t^{\prime}),{\mathbf{c}}_{i}(t^{\prime})]$ where ${\mathbf{c}}_{i}(t^{\prime})={\mathbf{c}}_{i}^{\prime}$, ${\bf{co}}_{i}(t+s)\in{\mathcal{N}}^{i}_{\text{rand}}(t)$ for all $s\in\{1,2,\dots,t^{\prime}-t\}$, and the starting configuration ${\mathbf{p}}(t)\in{\mathcal{N}}_{\text{rand}}(t)$ is within a platform-specific threshold $\epsilon$ from $o_{i}$, enabling grasping.
+Note that if $o_{i}$ is moved to ${\mathbf{c}}_{i}^{\prime}$, the set ${\mathcal{N}}^{i}_{\text{rand}}(t)$ may split into multiple components, each adjacent to ${\bf{co}}_{i}^{\prime}$.
+After the move, the robot may be located in any of these components; we randomly select one, denoted $\mathcal{N}$$^{\prime}$, and require ${\mathbf{p}}(t^{\prime})\in{\mathcal{N}}^{\prime}$.
+Existence of $\zeta$ can be verified using any motion planner (e.g., RRT [1]) [line 2].
+Under the considered holonomic dynamics (see Section II), the exact starting location ${\mathbf{p}}(t)$ of the robot does not affect the feasibility of transporting the object.
+If $\zeta$ exists, then we construct ${\mathbf{c}}(t^{\prime})$ that differs from ${\mathbf{c}}(t)$ only in the configuration of $o_{i}$ [line 2, Alg.
+2] and we set ${\mathcal{N}}(t^{\prime})={\mathcal{N}}^{\prime}$.
+Then, we construct the node ${\mathbf{q}}_{\text{new}}=[{\mathcal{N}}(t^{\prime}),{\mathbf{c}}(t^{\prime})]$ [line 2, Alg.
+2] and the sets of nodes and edges are updated as ${\mathcal{V}}={\mathcal{V}}\cup\{{\mathbf{q}}_{\text{new}}\}$ and ${\mathcal{E}}={\mathcal{E}}\cup({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ [lines 1, Alg.
 1].
-Also, once an edge $({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ is added to the tree, we store the corresponding manipulation $a(c_{i}(t),c_{i}^{\prime})$ that can enable it.
+[^1] We note that Alg.
+12 do not specify the exact time step $t^{\prime}$ at which the node ${\mathbf{q}}_{\text{new}}$ is reached.
+The time stamp is included only to indicate that ${\mathbf{q}}_{\text{new}}$ will be reached at some future step.
+The actual time step $t$ at which a tree node $\mathbf{q}$} is reached is determined by the robot’s motion path $\xi$ used to execute a plan $\tau$ that passes through $\mathbf{q}$}; computation of $\xi$ is discussed later.
+Also, once an edge $({\mathbf{q}}_{\text{rand}},{\mathbf{q}}_{\text{new}})$ is added to the tree, we store the corresponding manipulation $a({\mathbf{c}}_{i}(t),{\mathbf{c}}_{i}^{\prime})$ that can enable it; see Fig.
+2.
 We formalize this by iteratively constructing a function $A$ which returns the corresponding action for each edge in $\mathcal{E}$} [line 1, Alg.
 1].
-If a collision-free path $\boldsymbol{\zeta}$ does not exist or if $c_{i}^{\prime}$ is not a valid configuration, then we set $q_{\text{new}}=\varnothing$ and the iteration $m$ will terminate without adding any new nodes to the tree [lines 2-2, Alg.
+If a collision-free path $\zeta$ does not exist or if ${\mathbf{c}}_{i}^{\prime}$ is not a valid configuration, then we set $q_{\text{new}}=\varnothing$ and the iteration $m$ will terminate without adding any new nodes to the tree [lines 2-2, Alg.
 2].
@@
 1].
-Using $\tau_{{\mathbf{q}}}$ and the function $A$, we can extract the plan $\tau=\tau(0),\dots,\tau(H)$ modeling the sequence of manipulation actions that need to be applied to solve Problem 1 [line 1, Alg.
+Using $\tau_{{\mathbf{q}}}$ and the function $A$, we can extract the plan $\tau=\tau(0),\dots,\tau(H)$ [line 1, Alg.
 1].
-The node ${\mathbf{q}}_{g}$ can be selected randomly or based on any user-defined criterion such as the horizon $H$ of $\tau$ or the distance that needs to be traveled to execute $\tau$; see also Rem.
-III.1.
-Remark III.1 (Execution of $\tau$)
-Each action $\tau(k)=a(c_{i},c_{i}^{\prime})$ in $\tau$ involves moving an object $o_{i}$ from its current configuration $c_{i}$ to a new configuration $c_{i}^{\prime}$.
-Execution of this action should be in accordance with the corresponding path $\tau_{{\mathbf{q}}}$.
-This means that the robot should grasp $o_{i}$, while being in ${\mathcal{N}}$ as determined by $\tau_{{\mathbf{q}}}(k)=[{\mathcal{N}},{\mathbf{c}}]$, and drop-it off at $c_{i}^{\prime}$ while being in ${\mathcal{N}}^{\prime}$ determined by $\tau_{{\mathbf{q}}}(k+1)=[{\mathcal{N}}^{\prime},{\mathbf{c}}^{\prime}]$.
-According to line 2 in Alg.
-2, there must exist a collision-free path for the robot to execute this movement.
-Also, observe that, as in [22], our planner decouples the computation of the robot’s motion from the computation of obstacle movements.
-Specifically, it does not determine the exact grasp poses for the objects of interest but instead tracks the free-space component of the robot.
-Figure 4: Input prompt for GPT-4o and $K=1$ in Case Study V (Sec. IV).
+The node ${\mathbf{q}}_{g}$ can be selected randomly or based on any criterion (e.g., the horizon $H$ of $\tau$).
+Once $\tau=\tau(0),\dots,\tau(k),\dots\tau(H)$ is computed, a robot motion path $\xi$, defined as a sequence of robot configurations $\mathbf{p}$}, can be constructed sequentially across $k=0,\dots,H$.
+Each action $\tau(k)=a({\mathbf{c}}_{i},{\mathbf{c}}_{i}^{\prime})$ involves moving an object $o_{i}$ from its current configuration ${\mathbf{c}}_{i}$ to a new configuration ${\mathbf{c}}_{i}^{\prime}$.
+Execution of $\tau(k)$ follows the corresponding path $\tau_{{\mathbf{q}}}$: the robot moves towards $o_{i}$ to grasp it while in $\mathcal{N}$}, determined by $\tau_{{\mathbf{q}}}(k)=[{\mathcal{N}},{\mathbf{c}}]$, and then drops it at ${\mathbf{c}}_{i}^{\prime}$ while in $\mathcal{N}$$^{\prime}$, determined by $\tau_{{\mathbf{q}}}(k+1)=[{\mathcal{N}}^{\prime},{\mathbf{c}}^{\prime}]$.
+This path can be computed using any existing motion planner (e.g., RRT).
+A motion path executing $\tau(k)$ exists because (i) the robot is holonomic and (ii) by construction of Algorithm 2 (i.e., existence of collision-free paths $\zeta$ for each $\tau(k)$).
+Concatenating these motion paths for all $k$ yields the full robot motion path $\xi$ [line 1, Alg.
+1].
 III-B LLM-guided Sampling Strategy
-The sampling-based planner above relies on the distributions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ to generate ${\mathbf{q}}_{\text{new}}$.
-While these can be uniform, as e.g., in [22], such strategies often lead to slow plan synthesis, particularly as environmental complexity (e.g., the number of movable objects) increases.
-To accelerate the planning process, we propose non-uniform mass functions $f_{{\mathcal{V}}}$ and $f_{{\mathcal{O}}_{M}}$ that bias sampling toward directions where a feasible solution is more likely to be found.
-Our approach uses LLMs to identify these promising directions.
+The proposed sampling-based planner relies on the distributions $f_{{\mathcal{V}}}$, $f_{{\mathcal{O}}_{M}}$, and $f_{{\mathcal{C}}_{i}}$ to generate ${\mathbf{q}}_{\text{new}}$.
+While these can be uniform, as in [16], such strategies often result in slow plan synthesis, particularly as environmental complexity (e.g., the number of movable objects) increases.
+To accelerate planning, we introduce non-uniform mass functions $f_{{\mathcal{V}}}$ and $f_{{\mathcal{O}}_{M}}$ that bias sampling toward directions where feasible solutions are more likely to be found, while retaining a uniform density for $f_{{\mathcal{C}}_{i}}$.
+To realize this bias, our approach leverages LLMs to identify promising directions.
 First, we define $f_{{\mathcal{V}}}$ as follows:
@@
 where $p_{\text{rand}}\in[0,1)$ is a user-specified parameter and ${\mathcal{V}}_{v}\subseteq{\mathcal{V}}$; ${\mathcal{V}}_{v}$ will be defined later once $f_{{\mathcal{O}}_{M}}$ is defined.
-According to (1), a node ${\mathbf{q}}\in{\mathcal{V}}_{v}$ is selected to be the node ${\mathbf{q}}_{\text{rand}}$ with probability $p_{\text{rand}}$ while any node ${\mathbf{q}}\in{\mathcal{V}}$ is selected with probability $1-p_{\text{rand}}$.
-Notice that if $p_{\text{rand}}=0$, then $f_{{\mathcal{V}}}$ models uniform sampling.
-As $p_{\text{rand}}$ increases, the sampling strategy becomes more biased toward selecting nodes ${\mathbf{q}}\in{\mathcal{V}}_{v}$.
+Notice that if $p_{\text{rand}}=0$, then $f_{{\mathcal{V}}}$ models uniform sampling while as $p_{\text{rand}}$ increases, the sampling strategy becomes more biased toward selecting nodes ${\mathbf{q}}\in{\mathcal{V}}_{v}$.
 Second, we define the function $f_{{\mathcal{O}}_{M}}$ as follows:
-$f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})=\begin{cases}p_{\text{obs}}+\frac{1}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}=o_{\text{LLM}},\\ \frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~{}o_{i}\in\mathcal{O}_{M,\text{rand}},\end{cases}$ (2)
-where $p_{\text{obs}}\in[0,1)$ is a user-specified parameter and $o_{\text{LLM}}$ is manipulable obstacle, i.e., $o_{\text{LLM}}\in{\mathcal{O}}_{M}$, that is recommended by an LLM to relocate.
-A discussion on how to prompt an LLM to select $o_{\text{LLM}}$ is deferred to Section III-C.
-According to (2), the obstacle $o_{\text{LLM}}\in{\mathcal{O}}_{M}$ is selected for relocation with probability $p_{\text{obs}}$ while any obstacle $o_{i}\in{\mathcal{O}}_{M}$ is selected with probability $1-p_{\text{obs}}$.
-The probability of selecting $o_{i}\notin{\mathcal{O}}_{M}$ is 00.
-Observe that if $p_{\text{obs}}=0$, then $f_{{\mathcal{O}}_{M}}$ models uniform sampling.
-As $p_{\text{obs}}$ increases, the sampling strategy becomes more biased toward selecting $o_{\text{LLM}}$.
-Also, we note that it is possible that the LLM may fail to select a manipulable obstacle or it may generate nonsensical information.
-In such cases, we set $p_{\text{obs}}=0$ for that iteration so that a manipulable obstacle is selected randomly; see Section III-C.
-Next, we define the set ${\mathcal{V}}_{v}$ used for the construction of $f_{{\mathcal{V}}}$ in (1).
-First, we define a function $v:{\mathcal{V}}\rightarrow\mathbb{N}$ assigning a positive score to each node ${\mathbf{q}}\in{\mathcal{V}}$.
+$f_{{\mathcal{O}}_{M}}(o_{i}\mid o_{\text{LLM}},\mathcal{O}_{M,\text{rand}})=\begin{cases}p_{\text{obs}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|}},&\text{if}~o_{i}=o_{\text{LLM}},\\ \frac{1-p_{\text{obs}}}{|\mathcal{O}_{M,\text{rand}}|},&\text{if}~o_{i}\in\mathcal{O}_{M,\text{rand}},\end{cases}$ (2)
+where, $p_{\text{obs}}\in[0,1)$ is a user-specified parameter, and $o_{\text{LLM}}\in{\mathcal{O}}_{M}$ is a manipulable obstacle recommended by an LLM for relocation.
+How to prompt the LLM to select $o_{\text{LLM}}$ is discussed in Section III-C.
+According to (2), the probability of selecting $o_{i}\notin{\mathcal{O}}_{M}$ is 0.
+If $p_{\text{obs}}=0$, $f_{{\mathcal{O}}_{M}}$ reduces to uniform sampling; as $p_{\text{obs}}$ increases, sampling increasingly favors $o_{\text{LLM}}$; see Fig.
+3.
+If the LLM fails to select a manipulable obstacle or outputs invalid information, we set $p_{\text{obs}}=0$ for that iteration so a manipulable obstacle is chosen randomly; see Section III-C.
+Next, we define the set ${\mathcal{V}}_{v}$ used to construct $f_{{\mathcal{V}}}$ in (1).
+First, we define a function $v:{\mathcal{V}}\rightarrow\mathbb{N}$ assigning a positive score to each tree node $\mathbf{q}$}.
 This function is initialized as $v({\mathbf{q}}(0))=0$.
-Second, we show how the score of a newly added node ${\mathbf{q}}_{\text{new}}$ is generated.
+Second, we show how the score of a newly added node ${\mathbf{q}}_{\text{new}}$ is generated; see Fig.
+2.
 If ${\mathbf{q}}_{\text{new}}$ was constructed by relocating $o_{\text{LLM}}$, then we set $v({\mathbf{q}}_{\text{new}})=v({\mathbf{q}}_{\text{rand}})+1$.
@@
 Informally, $v({\mathbf{q}}_{\text{new}})$ counts how many nodes along the path connecting ${\mathbf{q}}_{\text{new}}$ to ${\mathbf{q}}(0)$ have been constructed based on LLM recommendations.
-Third, we construct the set ${\mathcal{V}}_{v}$ so that it collects the nodes with the highest score $v$, i.e., ${\mathcal{V}}_{v}=\{{\mathbf{q}}\in{\mathcal{V}}~{}|~{}v({\mathbf{q}})=\max_{{\mathbf{q}}^{\prime}\in{\mathcal{V}}}v({\mathbf{q}}^{\prime})\}$.
-Finally, we select a uniform density function for $f_{{\mathcal{C}}_{i}}$.
-Non-uniform density functions $f_{{\mathcal{C}}_{i}}$ can also be implemented similarly, e.g., by querying an LLM to select configurations $c_{i}^{\prime}$ for which collision-free paths exist; for simplicity, we abstain from this design choice.
+Third, we construct ${\mathcal{V}}_{v}$ so that it collects the nodes with the highest score, i.e., ${\mathcal{V}}_{v}=\{{\mathbf{q}}\in{\mathcal{V}}~|~v({\mathbf{q}})=\max_{{\mathbf{q}}^{\prime}\in{\mathcal{V}}}v({\mathbf{q}}^{\prime})\}$; see Fig.
+2.
 III-C Prompt Structure
-In this section, we describe how we query a pre-trained LLM to select a manipulable obstacle $o_{\text{LLM}}$.
-The input prompt to the LLM consists of three parts (see Fig.4): (1) Task description which provides information about the environment’s structure (i.e., the location of all movable and non-movable obstacles and the goal region) and specifies that the robot’s task is to remove obstacles so that the goal region is reachable; (2) Rules that the robot must follow to accomplish its task.
-Specifically, these rules state that the robot can only remove manipulable obstacles, should minimize the number of obstacles removed and should avoid picking manipulable obstacles that are blocked by other obstacles (see e.g., Case Study VI in Section IV); (3) Output format which defines how the LLM’s response should be formatted for compatibility with Alg.
-2.
-In particular, the LLM is required to generate $K\geq 1$ obstacles in the form $\text{obs}_{i}$ where $i$ is a unique identifier of each obstacle.
-Among the generated obstacles, one is selected randomly and denoted by $o_{\text{LLM}}$ in (2).
-Note that the LLM may make content and format errors.
-A content error occurs when the selected obstacle is not manipulable, and a format error occurs when the LLM’s output does not comply with the required format.
-If either of these errors occurs, we set $p_{\text{obs}}=0$.
+In this section, we describe how we prompt a pre-trained LLM to select a manipulable obstacle $o_{\text{LLM}}$ in a single query.
+The input prompt has three parts (see Fig.4): (1) Task description, which specifies the environment (locations of movable/non-movable obstacles and the goal region) and the robot’s task of removing obstacles to reach the goal; (2) Rules, which state that the robot may only remove manipulable obstacles, should minimize removals, and should avoid selecting blocked manipulable obstacles (see Case StudyVI in Section IV); and (3) Output format, which defines how the LLM’s response must be structured for compatibility with Alg.2.
+In particular, the LLM is asked to generate $K\geq 1$ obstacles (in a single query) in the form $\text{obs}i$, where $i$ uniquely identifies an obstacle.
+One of the generated obstacles is then selected at random and denoted $o_{\text{LLM}}$ in (2).
+If $o_{\text{LLM}}$ is not manipulable (content error) or the response does not follow the required structure (format error), we discard the LLM output and set $p_{\text{obs}}=0$ for that iteration.
+Figure 4: Input prompt for GPT-4o and $K=1$ in Case Study V (Sec. IV).
 III-D Probabilistic Completeness
-In what follows, we show that NAMO-LLM is probabilistically complete.
+Next, we show that NAMO-LLM is probabilistically complete.
 To formally state this result, we first introduce the notion of path clearance [16].
@@
 (a) The transitions from $\tau_{{\mathbf{q}}}(k-1)$ to $\tau_{{\mathbf{q}}}(k)$ and from $\tilde{\tau}_{{\mathbf{q}}}(k-1)$ to $\tilde{\tau}_{{\mathbf{q}}}(k)$ must involve relocating the same manipulable obstacle $o_{i}$, albeit potentially to different configurations.
-(b) Let $c_{i}^{k}$ and $\tilde{c}_{i}^{k}$ denote the configurations of that obstacle $o_{i}$ in $\tau_{{\mathbf{q}}}(k)$ and $\tilde{\tau}_{{\mathbf{q}}}(k)$, respectively.
-These configurations must satisfy $\lVert c_{i}^{k}-\tilde{c}_{i}^{k}\rVert\leq\epsilon$ for all $k$, for some $\epsilon>0$.
-In what follows, we refer to the satisfaction of conditions (i)–(ii) concisely as $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$.
-Theorem III.2
+(b) Let ${\mathbf{c}}_{i}^{k}$ and $\tilde{{\mathbf{c}}}_{i}^{k}$ denote the configurations of that obstacle $o_{i}$ in $\tau_{{\mathbf{q}}}(k)$ and $\tilde{\tau}_{{\mathbf{q}}}(k)$, respectively.
+These configurations must satisfy $\lVert{\mathbf{c}}_{i}^{k}-\tilde{{\mathbf{c}}}_{i}^{k}\rVert\leq\epsilon$ for all $k$, for some $\epsilon>0$.
+In what follows, we refer to satisfaction of conditions (i)–(ii) concisely as $\lVert\tau_{{\mathbf{q}}}(k)-\tilde{\tau}_{{\mathbf{q}}}(k)\rVert\leq\epsilon$.
+Theorem III.1
 Assume that there exists a feasible path $\tau_{{\mathbf{q}}}=\tau_{{\mathbf{q}}}(0),...,\tau_{{\mathbf{q}}}(H+1)$ with clearance $\epsilon>0$.
-If $p_{\text{rand}}\in[0,1)$, $p_{\text{obs}}\in[0,1)$, and the motion planner used to check existence of collision-free paths between obstacle configurations in line 2 of Alg.
-2 is complete, then the probability that NAMO-LLM will find a feasible path approaches $1$ as $m$ goes to $\infty$.
-Proof:
-To show this result, we need first to introduce the following notations and definitions.
+If $p_{\text{rand}}\in[0,1)$, $p_{\text{obs}}\in[0,1)$, and the motion planner used to check existence of collision-free paths in [line 2, Alg.
+2] is complete, then the probability that NAMO-LLM will find a feasible path approaches $1$ as $m$ goes to $\infty$.
+Proof: To show this result, we need first to introduce the following notations and definitions.
 Let ${\mathcal{G}}_{m}=\{{\mathcal{V}}_{m},{\mathcal{E}}_{m}\}$ denote the tree constructed by Algorithm 1 at the end of iteration $m$.
@@
 $2\sum_{1\leq i<j\leq m}\Bigl[\mathrm{E}(X_{i}X_{j})-p_{i}p_{j}\Bigr]<2\sum_{i=1}^{m}p_{i}=2\mathrm{E}(Y_{k})$
-Thus we get $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$ completing the proof of part (b). ∎
+Thus we get $\mathrm{Var}(Y_{m})<3\mathrm{E}(Y_{m})$ completing the proof of part (b).
 IV Experiments
-In this section, we present extensive numerical experiments to evaluate the efficiency of our method.
-Specifically, we demonstrate that our approach can effectively solve NAMO problems involving hundreds of movable obstacles and requiring the manipulation of multiple obstacles, significantly outperforming recent search-based methods [18] and sampling-based planners relying on uniform sampling [22] in terms of both runtime and plan quality.
-Hardware demonstrations of NAMO-LLM on ground robots are provided in [38].
-For these experiments, we use the GPT-4o and Gemini-1.5-flash models.
+In this section, we demonstrate that NAMO-LLM outperforms search-based methods [12], sampling-based planners [16], and LLM-based planners in terms of runtimes, plan quality, and task success rates.
+Our evaluation is performed across different LLMs and environmental setups.
+Hardware demonstrations of our method on ground robots are also provided; see also the supplemental material.
 All experiments were conducted on a computer with an Intel Core i5 2.4GHz processor and 8 GB of RAM.
@@
 IV-A Setting Up Comparative Experiments
-NAMO Baselines: We evaluate NAMO-LLM against two baselines: (i) RandomTree, a sampling-based approach from [22] that uses uniform sampling strategies, and (ii) NAMO-SA, a search-based method from [18].
-Notably, RandomTree can be recovered from our method by setting $p_{\text{rand}}=p_{\text{obs}}=0$.
-The approach in [18] selects obstacles for removal by computing the total distance from the robot to the obstacle and from the obstacle to the goal region using the $\text{A}^{*}$ algorithm.
-This is done for all manipulable objects, and the obstacle with the lowest total distance is selected.
-The chosen obstacle is then moved to a pre-defined storage room if a collision-free path towards that storage room exists.
-If no path to the storage room exists, the manipulable obstacle with the next lowest cost is selected; more details can be found in [18].
-In our implementation of [18], we model the storage room as an external region adjacent to a specified portion of the workspace boundary.
-Once an obstacle reaches this boundary segment, it is assumed to be instantaneously transferred to the storage room.
-On the other hand, our method designs custom drop-off locations for the obstacles during the sampling process.
-Note that this favors the baseline as our method may sample invalid drop-off locations (e.g., configurations lying in the obstacle space), resulting in sample rejection; see lines 2-2 in Alg.
+NAMO Baselines: We consider three baselines: (i) RandomTree, a sampling-based approach from [16] that employs uniform sampling; (ii) NAMO-SA, a search-based method from [12]; and (iii) an LLM-based planner that relies exclusively on LLMs.
+Among these, only (iii) incorporates LLMs.
+Comparisons with (i)–(ii) are reported in Sections IV-BIV-C, and comparisons with (iii) are presented in Section IV-D.
+RandomTree is recovered from our method by setting $p_{\text{rand}}=p_{\text{obs}}=0$, while the LLM-based planner corresponds to NAMO-LLM with $p_{\text{rand}}=p_{\text{obs}}=1$ and $K=1$.
+Unlike our method and RandomTree, the LLM-based planner lacks completeness guarantees (see Theorem III.1).
+RandomTree and NAMO-LLM run until either a feasible solution is found or a maximum of $m_{\text{max}}=30,000$ iterations is reached.
+In the LLM-based planner, the LLM is queried at every iteration to select an obstacle, which is placed at a random location.
+To ensure financially sustainable but fair comparisons, we cap the total number of LLM queries to $3H_{\text{min}}$, where $H_{\text{min}}$ is the minimum number of obstacles that need to be relocated to clear a path to the goal area.
+At each iteration, the LLM receives the same prompt as NAMO-LLM and outputs an obstacle.
+If the output has no format/content errors, the obstacle is placed randomly; otherwise, the counter is incremented and the LLM is queried again.
+The approach in [12] selects obstacles for removal by computing the total distance from the robot to the obstacle and from the obstacle to the goal region using $\text{A}^{*}$.
+The obstacle with the lowest cost is chosen and moved to a pre-defined storage room if a collision-free path exists; otherwise, the next lowest-cost obstacle is considered.
+In our implementation, the storage room is modeled as an external region adjacent to the workspace boundary, and once an obstacle reaches this boundary it is assumed to be instantaneously transferred.
+In contrast, our method designs custom drop-off locations during sampling.
+This favors the baseline, since our method may sample invalid drop-off locations leading to rejections; see lines 22 in Alg.
 2.
-Also in [18], once an obstacle is moved to the storage room, it is never considered for relocation again.
-This design choice favors that baseline, as the number of obstacles it needs to handle decreases over time—unlike our method.
-Thus, the maximum horizon $H$ of a plan designed by [18] is equal to the number $n$ of movable obstacles, while in ours it is possible that $H>n$ since our planner is allowed to reconfigure obstacles multiple times.
-We also note here that [18] includes re-planning strategies to handle unknown obstacles (i.e., obstacles with unknown locations and movability).
-Since our focus is on known environments, we exclude these capabilities.
+Moreover, in [12], once an obstacle is moved to the storage room it is never relocated, further favoring the baseline as the number $n$ of movable obstacles decreases over time—unlike in our method.
+Thus, the maximum horizon $H$ of a plan in [12] equals $n$, whereas in ours it is possible that $H>n$ since obstacles may be reconfigured multiple times.
+Re-planning strategies in [12] to handle unknown obstacles are neglected in our implementation as we focus on known environments.
 Finally, to ensure fair comparisons, our implementation of NAMO-LLM ([line 2, Alg.
 2]), [12], and [16] share the same RRT planner to check if there exists a collision-free path towards the drop-off location of an obstacle and the storage area.
-Robot and Obstacle Configuration Space: We consider a disk-shaped robot with radius $r=0.2$ (Case Studies I-IV) and $r=0.5$ (Case Studies V-VI) and configuration space ${\mathcal{P}}$ determined by the position of the center of the robot.
-We consider polygonal obstacles with configuration space ${\mathcal{C}}_{i}$ defined by their position and orientation.
-Evaluation Metrics: As evaluation metrics, we use (i) the runtime to compute the first feasible plan and (ii) the horizon $H$ of the designed plan.
-Runtimes and plan horizons $H$ for our method and the sampling-based baseline [22] are averaged over $100$ trials.
+Robot and Obstacle Configuration Space: We consider a disk-shaped robot of radius $r$ ($r=0.2$ in Cases I–IV, $r=0.5$ in Cases V–VI) with configuration space $\mathcal{P}$} given by the robot’s center position.
+Obstacles are polygonal, each with configuration space ${\mathcal{C}}_{i}$ defined by its position and orientation.
+Evaluation Metrics: As evaluation metrics we use (a) the runtime to compute the first feasible plan; (b) the horizon $H$ of the designed plan; and (c) task success rates.
+Task success rates, runtimes, and plan horizons $H$ for our method, the LLM-based planner, and [16] are averaged over $100$ runs.
+Task success rates are defined as the percentage of runs that result in a feasible manipulation plan.
+Note that NAMO-SA’s success rate is always $100\%$, since it has no randomized component and, in the worst case, moves all objects sequentially to a storage room as discussed earlier.
+RandomTree and NAMO-LLM will also have a $100\%$ success rate, for large enough $m_{\text{max}}$, due to their completeness properties.
+Our reported times include all computational components of Alg.
+1–2 including e.g., the runtime for LLM inference.
+Figure 5: Initial configurations of the obstacles (red polygons) and robot for Case Studies I-IV.
+The green path connects the initial robot state to the first obstacle to be repositioned for clearing a path to the goal area ${\mathcal{P}}_{g}$.
+IV-B Comparisons with Search- and Sampling-Based Baselines
 […]
-IV-B Comparative Experiments against NAMO Baselines
-In what follows, we present six case studies that differ in their complexity in terms of the total number $n$ of movable obstacles, the initial configuration of the obstacles, and the minimum horizon, denoted by $H_{\text{min}}$, of a plan $\tau$ required to reach ${\mathcal{P}}_{g}$.
-In our method, NAMO-LLM$(p_{\text{rand}},p_{\text{obs}})$, we employ distinct parameter settings for $p_{\text{rand}}$ and $p_{\text{obs}}$.
-The following results are obtained using the pre-trained GPT 4o model.
 Method $|{\mathcal{V}}|$ Time (secs) $H$
@@
 NAMO-LLM(0.8, 0.8) 3.21 4.25 2.67
-[…]
+TABLE I: Case Study I: Average tree size $|{\mathcal{V}}|$, runtimes, and horizon $H$.
 Method $|{\mathcal{V}}|$ Time (secs) $H$
-[…]
+Case II NAMO-SA [12] N/A 876.36 65
 RandomTree [16] 34.90 42.78 4.57
-5
 NAMO-LLM(0.2, 0.8) 3.86 9.99 2.75
 NAMO-LLM(0.8, 0.8) 4.47 7.80 3.04
-[…]
+Case III NAMO-SA [12] N/A 151.23 26
 RandomTree [16] 92.72 71.32 5.72
-5
 NAMO-LLM(0.2, 0.8) 6.33 8.52 3.70
 NAMO-LLM(0.8, 0.8) 4.37 5.91 3.74
-[…]
+Case IV NAMO-SA [12] N/A 877.97 65
 RandomTree [16] 290.83 408.68 7.41
-5
 NAMO-LLM(0.2, 0.8) 8.23 19.31 4.16
 NAMO-LLM(0.8, 0.8) 5.70 17.12 4.5
-[…]
-Case Study I ($n=50,H_{\text{min}}=1$): In this case study, we consider a $50\times 50$ environment with $n=50$ movable obstacles.
-The initial configurations of the obstacles and the robot, as well as the desired robot configuration, are depicted in Fig.
+TABLE II: Case Study II-IV: Average tree size, runtimes, and horizon.
+Case Study I ($n=50,H_{\text{min}}=1$): We consider a $50\times 50$ environment with $n=50$ movable obstacles admitting a feasible plan with $H_{\text{min}}=1$; see Fig.
 5.
-The environment is designed such that a plan with horizon $H_{\text{min}}=1$ exists to reach ${\mathcal{P}}_{g}$, meaning that the minimum number of obstacles that must be reconfigured to reach ${\mathcal{P}}_{g}$ is $1$.
-We test NAMO-LLM for various values of the parameters $p_{\text{rand}}$, $p_{\text{obs}}$, and $K$.
-The comparative results are summarized in Table I.
+We compared NAMO-LLM for various values of $p_{\text{rand}}$, $p_{\text{obs}}$, and $K$ against the baselines; see Table I.
 Observe that our method significantly outperforms both baselines in terms of both the runtime required to compute the first feasible plan and the horizon $H$ of that plan.
-For instance, the runtimes of the search- and sampling-based methods are $99.05$ and $14.94$ seconds, respectively, while our method can compute the first feasible plan significantly faster; e.g., the average runtime for $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$ is $2.88$ seconds.
-The average horizon of the plan computed by our method is consistently less than $4$, which is comparable to that of the sampling-based baseline, while the horizon of the search-based one is $27$.
-We note that the performance of our method does not change significantly with respect to $K\in\{1,2,3\}$.
-We attribute this to the fact that there are multiple candidate obstacles near ${\mathcal{P}}_{g}$ that can be removed to get access to it and that the employed LLM is capable of identifying them.
-Case Study II ($n=100,H_{\text{min}}=1$): In this case study, we consider a $65\times 65$ environment containing $n=100$ movable obstacles, where ${\mathcal{P}}_{g}$ can be reached by removing at least one obstacle; see Fig 5.
-The comparative results are shown in Table II.
-Due to space limitations, we present results for our method only with $K=1$.
-Observe that the performance gap between our method and the baselines increases as $n$ grows.
-For instance, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, our method finds the first feasible plan in $7.80$ seconds with an average horizon of $H=3.04$.
-In contrast, [22] takes $42.78$ seconds to compute a plan with a slightly longer horizon ($H=4.57$), while [18] requires $876.36$ seconds and produces a plan with $H=65$.
-Case Study III ($n=50,H_{\text{min}}=2$): Here we consider a $50\times 50$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5.
-The results in Table II show that the performance gap between our method and the baselines further increases as the minimum required horizon $H_{\text{min}}$ grows.
-Case Study IV ($n=100,H_{\text{min}}=2$): This case study examines a more complex environment with dimensions $65\times 65$ and $n=100$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least two obstacles; see Fig 5.
-The results are summarized in Table II, showing that the performance gap further increases as both $n$ and the minimum horizon $H_{\text{min}}$ increase.
-For instance, our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, finds the first feasible plan in $17.12$ seconds with an average horizon of $H=4.5$.
-In contrast, the sampling-based approach takes $408.68$ seconds to compute a plan with a longer horizon of $7.41$, while the search-based method requires $877.97$ seconds and produces a plan with $H=65$.
-[…]
-Case Study V ($n=50,H_{\text{min}}=6$): In this case study, we consider a $40\times 40$ environment with $n=50$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires removing at least six obstacles (see Fig.
-6).
-This scenario is significantly more complex due to the larger number of obstacles that need to be relocated.
-This complexity is also reflected in the increased runtimes reported in Table III, compared to the previous case studies.
-Our method, when configured with $K=1$, $p_{\text{rand}}=0.8$, and $p_{\text{obs}}=0.8$, computed the first feasible plan in $48.22$ seconds, with an average horizon of $H=26.48$.
-In contrast, [22] failed to find a feasible plan within $6$ hours while [18] can identify a feasible plan within $141.91$ seconds with $H=42$.
-Case Study VI ($n=10,H_{\text{min}}=3$): Finally, we consider a more complex scenario to illustrate the utility of parameter $K$, where—unlike previous case studies—some manipulable obstacles are blocked by others and cannot be relocated until the blocking ones are moved.
-Specifically, we consider a $20\times 20$ environment with $n=10$ movable obstacles, where reaching ${\mathcal{P}}_{g}$ requires relocating the manipulable obstacle $o_{1}$.
-However, unlike Case Studies I-IV, there do not exist collision-free paths to any possible configuration $c_{1}^{\prime}$ unless both $o_{8}$ and $o_{2}$ are moved first; see Fig 6.
-Thus, in this case, at least $3$ obstacles must be relocated to reach ${\mathcal{P}}_{g}$.
-The comparative results are shown in Table IV.
-Unlike Case Study I, the performance of our method improves significantly for $K\in\{2,3\}$ compared to $K=1$.
-This occurs because, with $K=1$, the LLM often prioritizes removing $o_{1}$ first (i.e., $o_{\text{LLM}}=o_{1}$ in (2)), even though it cannot be moved due to being blocked by $o_{8}$ and $o_{2}$.
-Consequently, any configuration $c_{1}^{\prime}$ sampled for $o_{1}$ is rejected by line 2 in Alg.
-2, slowing down the overall algorithm.
-However, when $K>1$, the LLM generates a set of obstacles that need to be removed to access ${\mathcal{P}}_{g}$, which may include $o_{2}$ and/or $o_{8}$.
-Randomly selecting an obstacle from this set helps the algorithm avoid getting stuck due to persistent incorrect LLM recommendations and, consequently, improves performance.
-Also, observe that our method outperforms [22], which relies on random exploration, in terms of runtime—even when configured with $K=1$.
-We attribute this to the fact that the LLM occasionally provides correct guidance (i.e., recommending the removal of $o_{2}$ and $o_{8}$ before $o_{1}$) after some redundant actions have been applied (e.g., relocating $o_{3}$ first), as these actions influence the input prompt to the LLM.
-Overall, our method outperforms both baselines when $K$ is $2$ or $3$.
-Observe in Table IV that [18] eventually moved all obstacles to the storage area leading to plans with $H=10$.
-Summary of Comparisons: Overall, the performance gap between NAMO-LLM and the baselines increases as $n$ or $H_{\text{min}}$ grows.
-In general, we observed that our method tends to achieve its lowest runtime for computing the first feasible plan when configured with $p_{\text{rand}}=p_{\text{obs}}=0.8$ and $K=1$; if manipulable obstacles are not blocked by other obstacles; see e.g., Case Studies I–V.
-We attribute this to the LLM’s ability to correctly reason about which obstacle to remove in these scenarios.
-However, when $p_{\text{rand}}=0.2$ and $p_{\text{obs}}=0.8$, the generated plans tend to have a shorter horizon $H$.
-This occurs because a smaller $p_{\text{rand}}$ enables the algorithm to select tree nodes for expansion more uniformly, leading to better-quality first plans, albeit with slightly higher runtimes.
-We also observed that pre-trained LLMs may occasionally make infeasible recommendations regarding which manipulable obstacle to relocate, especially when these obstacles are blocked by others and cannot be repositioned until the blocking ones are moved first; see e.g., Case Study VI.
-In such cases, setting $K>1$ is beneficial, as it prevents the algorithm from getting stuck due to poor LLM recommendations.
-Finally, we note that these results depend on the selected LLM.
+For instance, NAMO-LLM tends to be 30x and 5x faster than [16] and [12], respectively.
+Observe that the performance of our method does not change significantly with respect to $K\in\{1,2,3\}$.
+We attribute this to the fact that there are multiple candidate obstacles near the goal area ${\mathcal{P}}_{g}$ that can be removed to get access to it and that the employed LLM is capable of identifying them.
+In Case Studies II-V, we present results for our method with $K=1$ combined with $p_{\text{rand}}=p_{\text{obs}}=0.8$, and $p_{\text{rand}}=0.2$, $p_{\text{obs}}=0.8$.
+In Case Study II ($n=100,H_{\text{min}}=1$), Case Study III ($n=50,H_{\text{min}}=2$) and Case Study IV ($n=100,H_{\text{min}}=2$), we similarly observe that the performance gap between our method and the baselines further increases as the minimum required horizon $H_{\text{min}}$ grows and/or $n$ increase.
+The environments associated with these case studies can be found in Fig 5 while the results are shown in Table II.
+Case Study V ($n=50,H_{\text{min}}=6$): We consider a $40\times 40$ environment; see Fig.
+6.
+This scenario is significantly more complex due to the larger number of obstacles that need to be relocated; this is also reflected in the increased runtimes reported in Table III, compared to the previous case studies.
+Our method, when configured with $K=1$, $p_{\text{rand}}=p_{\text{obs}}=0.8$ computed the first feasible plan in $48.22$ secs, with an average horizon of $H=26.48$.
+In contrast, [16] failed to find a feasible plan within $6$ hours while [12] can identify a feasible plan within $141.91$ secs with $H=42$.
+Figure 6: Case Study V & VI: Initial configurations of the obstacles (red polygons). The green area corresponds to ${\mathcal{P}}_{g}$.
+Case Study VI ($n=10,H_{\text{min}}=3$): Finally, we consider a more complex scenario to illustrate the utility of parameter $K$.
+Unlike previous case studies, here some manipulable obstacles are blocked by others.
+Specifically, we consider a $20\times 20$ environment with $n=10$ movable obstacles, where reaching $\mathcal{P}_{g}$ requires relocating $o_{1}$, but no collision-free path to any ${\mathbf{c}}_{1}^{\prime}$ exists until $o_{8}$ and $o_{2}$ are moved; see Fig.
+6.
+Table IV shows that, unlike Case I, performance improves significantly for $K\in\{2,3\}$ compared to $K=1$.
+This occurs because because with $K=1$ the LLM often prioritizes $o_{1}$, which is blocked by $o_{8}$ and $o_{2}$; any ${\mathbf{c}}_{1}^{\prime}$ sampled for $o_{1}$ is rejected at line 2 in Alg.
+2, slowing progress.
+However, for $K>1$, the LLM outputs a set of candidate obstacles (often including $o_{2}$ and/or $o_{8}$), and random selection from this set helps avoid persistent incorrect recommendations.
+Even with $K=1$, our method outperforms [16] in runtime because the LLM occasionally recommends $o_{2}$ and $o_{8}$ after some exploratory moves (e.g., relocating $o_{3}$), which alters its prompt.
+Overall, our method surpasses both baselines for $K=2,3$; note in Table IV that [12] moved all obstacles to storage since $H=10$.
+Summary: The performance gap between NAMO-LLM and the baselines increases as $n$ or $H_{\text{min}}$ grows.
+Our method achieves its lowest runtime for computing the first feasible plan when $p_{\text{rand}}=p_{\text{obs}}=0.8$ and $K=1$, provided manipulable obstacles are not blocked (Cases I–V).
+This is due to the LLM’s ability to identify the correct obstacle to remove in these scenarios.
+When $p_{\text{rand}}=0.2$ and $p_{\text{obs}}=0.8$, plans typically have shorter horizons, since more uniform node expansion leads to higher-quality first plans, albeit with slightly longer runtimes.
+We also observed that LLMs may produce infeasible recommendations when manipulable obstacles are blocked (e.g., Case VI).
+In such cases, setting $K>1$ prevents the algorithm from stalling on poor LLM predictions.
+Finally, across all case studies, $2.77\%$ of LLM outputs are invalid (format or content errors) on average, while $\sim 70\%$ of which are due to content errors; these are handled via uniform sampling as discussed in Section III-C.
 Method $|{\mathcal{V}}|$ Time (secs) $H$
@@
 NAMO-LLM(0.8, 0.8) 34.78 48.22 26.48
-[…]
+TABLE III: Case Study V: Average tree size, runtimes, and horizon.
 Method $|{\mathcal{V}}|$ Time (secs) $H$
@@
 NAMO-LLM(0.8, 0.8) 12.65 42.47 9.9
-[…]
+TABLE IV: Case Study VI: Average tree size, runtimes, and horizon.
 Method $|{\mathcal{V}}|$ Time (secs) $H$
-[…]
+Case IV NAMO-SA [12] N/A 877.97 65
 RandomTree [16] 290.83 408.68 7.41
-NAMO-LLM(0.2, 0.8) 84.35 233.94 15.8
-NAMO-LLM(0.8, 0.8) 24.75 104.30 17.2
 GPT 4o 5.70 17.12 4.16
-[…]
+MiniGPT4o(0.2, 0.8) 19.98 81.45 5.62
+MiniGPT4o(0.8, 0.8) 10.18 59.72 6.96
+Gemini(0.2, 0.8) 84.35 233.94 15.8
+Gemini(0.8, 0.8) 24.75 104.30 17.2
+Case V NAMO-SA [12] N/A 141.91 42
 RandomTree [16] over 30k over 22k N/A
-NAMO-LLM(0.2, 0.8) 27.70 21.03 9.6
-NAMO-LLM(0.8, 0.8) 15.15 14.24 12.15
 GPT 4o 34.78 48.22 19.65
-TABLE V: Effect of LLM Model on average tree size, runtimes, and horizon. The line ‘GPT 4o’ reports the best results of our method when paired with GPT 4o.
-IV-C Comparative Results with an Alternative LLM
-In this section, we demonstrate the performance of our method when paired with the Gemini-1.5-flash model on Case Studies IV ($n=100$, $H_{\text{min}}=2$) and V ($n=50$, $H_{\text{min}}=6$).
-The comparative results are shown in Table V.
-In Case Study IV, our method outperforms both baselines in terms of runtime but produces longer plans than the sampling-based baseline (while still shorter than those of the search-based one).
-In Case Study V, however, our approach achieves better performance than both baselines in terms of both runtime and plan horizon $H$.
-Overall, we observe that switching from GPT-4o to Gemini-1.5-flash tends to improve the performance of our algorithm in scenarios where manipulable obstacles are blocked by other obstacles (see Case Study V) and deteriorate its performance in cases with large number of movable obstacles (see Case Study IV).
-Nevertheless, our method still remains more scalable than the baselines.
+MiniGPT4o(0.2, 0.8) 52.1 67.61 13.23
+MiniGPT4o(0.8, 0.8) 20.3 27.13 13.57
+Gemini(0.2, 0.8) 27.70 21.03 9.6
+Gemini(0.8, 0.8) 15.15 14.24 12.15
+TABLE V: Performance comparison of different models in Cases IV and V, showing average tree size, runtimes, and horizon.
+The line ‘GPT 4o’ reports the best results of our method when paired with GPT 4o.
+Lines of the form ‘LLM($p_{\text{rand}},p_{\text{obs}}$)’ refer to our method when coupled with the respective LLM.
+IV-C Comparative Results with Alternative LLMs & VLMs
+We evaluate NAMO-LLM with GPT-4o, GPT-4o-mini, and Gemini-1.5-Flash on Cases IV and V; see the Table V).
+Switching among these models affects performance, but our method remains more scalable than the baselines.
+In Case Study IV, it outperforms both baselines in runtime for all tested LLMs, though with the Gemini model it produces longer plans than [16] (while still shorter than those of [12]).
+In Case Study V, our method outperforms both baselines in both metrics for every LLM tested.
+We also observed that the performance of our method deteriorates when it is paired with VLMs (GPT 4o).
+For instance, our average runtime and horizon on Case IV with $p_{\text{rand}}=p_{\text{obs}}=0.8$ is $252.21$ secs and $32.8$, respectively, while it fails to address Case V within $1,000$ secs.
+The VLM is prompted with an overhead image of the environment, replacing the textual description of the environment.
+We attribute this result to potentially limited VLM capabilities in object recognition and counting tasks.
+IV-D Comparisons with LLM-based Planners
+We compare NAMO-LLM against the LLM-based planner from Section IV-A, using GPT-4o.
+The baseline’s success rates in Cases I–VI are $86\%$, $92\%$, $78\%$, $90\%$, $54\%$, and $0\%$, respectively, while NAMO-LLM achieves $100\%$ in all.
+GPT-4o alone is competitive in simpler tasks (Cases I–IV) but drops sharply in complex ones (Cases V–VI).
+[^2] Other LLMs showed the same trend; e.g., GPT-4o-mini achieved 81%, 70%, 92%, 82%, 55%, and 0% for Cases I–VI, respectively.
+For Cases I–IV, both methods produce plans with comparable horizons.
+As complexity increases (Cases V–VI), NAMO-LLM outperforms the baseline.
+Across Cases I–VI, the best (average) horizons of NAMO-LLM are $2.5$, $2.75$, $3.70$, $4.16$, $19.65$, and $8.11$, compared to $2.68$, $2.7$, $3.4$, $3.75$, $52.88$, and N/A for the baseline.
+For runtimes, the baseline is slightly faster in simpler tasks but fails on more complex ones (Case V): NAMO-LLM achieves $2.88$, $7.80$, $5.91$, $17.12$, and $31.05$ secs across Cases I–V, compared to $2.73$, $6.47$, $5.37$, $13.17$, and N/A for the baseline.
+Overall, these comparisons show that our algorithm consistently outperforms the baseline in terms of task success rates, particularly in complex scenarios (e.g., Case V).
+The main advantage of the baseline is that, if it produces a plan, the plan may be found slightly faster especially in simpler tasks (Cases I–V).
+IV-E Hardware Validation Example
+To demonstrate the practical feasibility of transferring NAMO-LLM plans from simulation to real-world execution, we validated the method on a TurtleBot3 Waffle Pi with a 4-DOF OpenManipulator-X, operating in a 3D Mini-City platform; see Fig.
+1 and the supplemental material.
+The robot performed state estimation, navigation, and manipulation using off-the-shelf ROS packages.
+The experiment involved $n=19$ movable obstacles (traffic cones, construction blocks) and several fixed objects (buildings, pedestrian islands), with a minimal horizon $H_{\min}=2$.
+To facilitate computation of free-space components ${\mathcal{N}}(t)$ we over-approximated the robot as a disk of radius $r=15cm$ and the obstacles as polygons.
+Our algorithm found the first feasible path in $1.14$ secs on average and the optimal $H_{\min}$ plan in $7.32$ secs.
 V Conclusion
-In this paper, we proposed NAMO-LLM, a probabilistically complete sampling-based planner for NAMO problems.
-Through comparative experiments, we demonstrated its scalability and efficiency in highly cluttered environments.
-Future work will focus on extending NAMO-LLM to settings with obstacles of unknown location and movability.
+This paper proposed a probabilistically complete planner for NAMO problems.
+Comparative experiments verified its efficiency in cluttered environments.
+Future work will focus on extending to settings with unknown obstacle location and movability, and developing variants that avoid explicit C-space computation to better accommodate complex geometries and high-dimensional configuration spaces.